### v1

# Learning Graph Quantized Tokenizers for Transformers

Abstract
Transformers serve as the backbone architectures of Foundational Models, where a domain-specific tokenizer helps them adapt to various domains. Graph Transformers (GTs) have recently emerged as a leading model in geometric deep learning, outperforming Graph Neural Networks (GNNs) in various graph learning tasks. However, the development of tokenizers for graphs has lagged behind other modalities, with existing approaches relying on heuristics or GNNs co-trained with Transformers. To address this, we introduce GQT (Graph Quantized Tokenizer), which decouples tokenizer training from Transformer training by leveraging multi-task graph self-supervised learning, yielding robust and generalizable graph tokens. Furthermore, the GQT utilizes Residual Vector Quantization (RVQ) to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities. By combining the GQT with token modulation, a Transformer encoder achieves state-of-the-art performance on 16 out of 18 benchmarks, including large-scale homophilic and heterophilic datasets.
1 Introduction
Following the success of Transformers (Vaswani et al., 2017) in natural language processing (Devlin et al., 2019; Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021), Graph Transformers (GTs) (Dwivedi & Bresson, 2020; Ying et al., 2021a; Rampášek et al., 2022; Ma et al., 2023; Shirzad et al., 2023; Kong et al., 2023b; Chen et al., 2023; Wu et al., 2022b) have emerged as strong models in geometric deep learning. Unlike message-passing Graph Neural Networks (GNNs), which rely on strong locality inductive biases (Battaglia et al., 2018; Veličković et al., 2018; Hou et al., 2020; Hamilton et al., 2017a; Kipf & Welling, 2017), GTs are inherently more expressive due to their ability to capture long-range interactions between nodes (Ma et al., 2023). This is particularly beneficial in heterophilous settings where local alignment does not hold (Fu et al., 2024). GTs possess an expressive power at least equivalent to the 2-Weisfeiler-Lehman (WL) isomorphism test (Kim et al., 2022), which is sufficient for most real-world tasks (Zopf, 2022). This surpasses the expressive power of message-passing GNNs, which are limited to the 1-WL test (Ying et al., 2021a). Furthermore, a Transformer with sufficient attention heads can match or exceed the expressive power of a second-order invariant graph network, outperforming message-passing GNNs (Kim et al., 2022). However, both GNNs and Transformers are susceptible to over-smoothing (Li et al., 2018; Zhou et al., 2021; Dovonon et al., 2024).
GTs require consideration of both graph structure and features, as nodes with identical features will otherwise be projected into the same representation regardless of their surrounding structures (Hoang et al., 2024). There are three general approaches to address this limitation (Hoang et al., 2024): (1) node feature modulation, which involves injecting structural information into the node features; (2) context node sampling, where a sampling strategy is used to construct a sequence over the neighbor nodes; and (3) modifying the architecture of a vanilla Transformer to directly incorporate structural biases. Given that Transformers are universal approximators of sequence-to-sequence functions (Yun et al., 2020) and considering the rapid developments in efficient implementation of multi-head attention (MHA) module (Dao et al., 2022a; Liu et al., 2024), which enables longer context sizes of up to million-scale tokens (Reid et al., 2024), a well-designed graph tokenizer can allow a vanilla Transformer model to efficiently process even large-scale graphs. Recent studies on applying Large Language Models (LLMs) to graph-related tasks have found that representing graphs through textual descriptions can lead to surprisingly strong performance gains that surpass those of GNNs, suggesting that vanilla Transformers are indeed capable of effectively learning graph structures (Ye et al., 2024; He et al., 2024). Nonetheless, LLMs are not inference-efficient, and hence our goal in this paper is to devise a lightweight and efficient graph tokenization strategy that enables vanilla Transformer encoders to learn graph structures effectively, without relying on LLMs.
Tokenizers typically employ self-supervised objectives to abstract data into a sequence of discrete tokens, enabling Transformers to learn representations across various modalities as a unified stream of data. This discretization is achieved through vector quantization techniques (Van Den Oord et al., 2017; Lee et al., 2022), which offer several benefits, including: (1) significantly reduced memory requirements, (2) improved inference efficiency, (3) allowing Transformers to focus on long-range dependencies rather than local information, and (4) the capacity to learn more high-level representations due to a compact latent space (Yuan et al., 2021; Yu et al., 2022). These advantages are particularly important in auto-regressive generative modeling, where quantized tokens allow Transformers to generate high-quality outputs in multiple modalities (Dubey et al., 2024; Lee et al., 2022; Dhariwal et al., 2020; Ramesh et al., 2021; Team, 2024). Despite its importance in other domains, tokenization remains under-explored for graph-structured data. To address this limitation, we propose the Graph Quantized Tokenizer (GQT), a novel approach that learns a hierarchical sequence of tokens over graphs using self-supervised objectives tailored to graph-structured data. More specifically, our contributions are as follows:
  • We propose a graph tokenizer that utilizes multi-task graph self-supervised objectives to train a graph encoder, enabling it to fully capture local interactions and allowing the Transformer to focus on long-range dependencies.
  • Our approach adapts Residual Vector Quantization (RVQ) within the graph tokenizer to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities.
  • We introduce a novel combination of semantic edges and random walks to facilitate the Transformer’s access to long-range interactions, and employ hierarchical encoding and gating mechanisms to modulate the tokens and provide informative representations to the Transformer.
  • Through extensive experiments on both homophilic and heterophilic datasets, including large-scale benchmarks, we demonstrate that our proposed tokenizer enables a Transformer encoder to achieve state-of-the-art performance on 16 out of 18 benchmarks while substantially reducing the memory footprint of the embeddings.
2 Related Works
Graph Transformers (GTs) have shown promising performance on various graph learning tasks, surpassing GNNs on many benchmarks. GTs can be broadly categorized into two directions (Hoang et al., 2024; Müller et al., 2024): (1) modifying the vanilla Transformer architecture to incorporate structural inductive biases, or (2) encoding the input graph to make it compatible with the vanilla Transformer design. Early examples of the first approach include Graph Attention Network (Veličković et al., 2018), which uses an attention module to compute pairwise node attentions and masks the attention matrix based on connectivity information. Subsequent works have replaced the scaled-dot attention module with various structure-aware sparse attention modules (Rampášek et al., 2022; Bo et al., 2023; Ying et al., 2021a; Deng et al., 2024; Wu et al., 2023b; Liu et al., 2023a; Chen et al., 2022; Dwivedi & Bresson, 2020; Shirzad et al., 2023; Ma et al., 2023). Examples of the second approach include Graph Memory Network (Khasahmadi et al., 2020), which passes non-linear projections of node features and structural encoding to a Transformer-like model. Structural encoding methods, such as Laplacian eigenvectors or Random walk-based encoding (Dwivedi et al., 2022; Ma et al., 2023; Cantürk et al., 2024), allow injecting structural information directly into the node features. Another approach involves using GNNs to encode local structure along with node features, followed by passing the representation to vanilla Transformers to capture long-range dependencies. (Rong et al., 2020; Wu et al., 2021; Chen et al., 2023; 2022). Recent studies leverage LLMs, where graphs are represented through natural language expressions, and an LLM performs graph-related tasks through in-context learning, instruction tuning, or soft prompts (Fatemi et al., 2024; Ye et al., 2024; He et al., 2024). For a detailed survey on GTs, see (Müller et al., 2024; Hoang et al., 2024).
Graph Tokenization provides GTs with rich node tokens that encapsulate both structural and semantic information. Various approaches have been proposed to define these node tokens. TokenGT (Kim et al., 2022) treats nodes and edges as independent tokens defined by their features, type identifiers, and structural encodings. NAGphormer (Chen et al., 2023) represents each node as a set of $L$ tokens, where the $l^{th}$ token is the representation of the node from $l^{th}$ hop aggregation. In contrast, GraphiT (Mialon et al., 2021) defines a node token as the concatenation of its feature and representation from a graph convolutional kernel network (GCKN). VCR-Graphormer (Fu et al., 2024) expands the notion of node tokens to include sequences comprising the node feature and features of semantically and community-related neighboring nodes. SGT (Liu et al., 2023b) is a non-parametric tokenizer designed for molecular tasks, which simplifies the tokenization process to a non-parametric graph operator without non-linearity, unlike motif-based tokenizers (Zhang et al., 2021; Jin et al., 2018) or GNN pre-training methods (Xia et al., 2023). NodePiece (Galkin et al., 2022) is a knowledge-graph tokenizer that represents a target node as a hash of its top-k closest anchors, their distances, and relational context. For a more detailed review see Müller et al. (2024). While vector quantization (VQ) (Van Den Oord et al., 2017; Lee et al., 2022) has been explored in other modalities, its application in graph learning is limited. Notable exceptions include VQGraph (Yang et al., 2024), which employs VQ for graph distillation, and NID (Luo et al., 2024a), which uses VQ to learn discrete node IDs for downstream prediction tasks.
3 Preliminaries
Messag-Passing GNNs. Let $\mathcal{G}$ denote the space of graphs. A graph $g$ within this space is defined as $g=\left(\mathcal{V},\mathcal{E},\textbf{X},\textbf{E}\right)$ where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. $\textbf{X}\in\mathbb{R}^{\left|\mathcal{V}\right|\times d_{x}}$ represents the node features of dimension $d_{x}$, and $\textbf{E}\in\mathbb{R}^{\left|\mathcal{V}\right|\times\left|\mathcal{V}\right|\times d_{e}}$ represents the edge features of dimension $d_{e}$. A message-passing GNN takes $g$ as input and learns representations $h^{l}_{v}$ for $v\in\mathcal{V}$ ($h^{0}_{v}=x_{v}$) in each layer $l$ as follows (Gilmer et al., 2017):
$h^{l}_{v}=f^{l}_{\theta}\left(h^{l-1}_{v},g^{l}_{\phi}\left(\left\{\left(h^{l-1}_{v},h^{l-1}_{u},e_{uv}\right)|u\in\mathcal{N}_{i}(v)\right\}\right)\right)$ (1)
where $f_{\theta}$ and $g_{\phi}$ are known as update (combine) and message (aggregate) functions, respectively, and $\mathcal{N}_{i}(v)$ denotes the set of immediate neighbors of the node $v$. With this representation, we can perform various tasks, including node classification as $\text{MLP}\left(h_{v}\right)$, edge prediction as $\text{MLP}\left(h_{u}\odot h_{v}\right)$, or graph classification as $\text{MLP}\left(\mathcal{R}\left(\left\{h_{u}|u\in\mathcal{V}\right\}\right)\right)$ where $\mathcal{R}$ is a pooling (readout) function.
Graph Transformers, on the other hand, first use a tokenizer $T_{v}=\mathcal{T}_{\psi}\left(\mathcal{N}(v)\right)$ to map each node $v\in\mathcal{V}$ into a sequence of tokens $T_{v}$ by considering some notion of neighborhood $\mathcal{N}$. The simplest design is when $\mathcal{N}$ is zero-hop neighborhood (i.e., the node itself) and $\mathcal{T}_{\psi}$ is a node feature lookup function. The neighborhood $\mathcal{N}$ can be extended to include an ego network (Zhao et al., 2021) or top-k Random Walk based neighbors (Fu et al., 2024), and $\mathcal{T}_{\psi}$ can be enhanced to representations from a GNN (Chen et al., 2023). Once the tokens are computed, along with a node positional encoding function (PE), we can define the input to a Transformer as $h^{0}_{v}=\left[T_{v}||\text{PE}\left(v\right)\right]$ and compute the representation in each layer $l$ of a vanilla Transformer encoder as follows:
$h^{l}_{v}$ $=\text{LN}\left(\text{MHA}\left(\text{LN}\left(h^{l-1}_{v}\right)\right)+h^{l-1}_{v}\right)$ (2)
$h^{l}_{v}$ $=h^{l}_{v}+\text{MLP}\left(h^{l}_{v}\right)$ (3)
[…]
Vector Quantization projects embeddings $\textbf{X}\in\mathbb{R}^{n\times d_{x}}$ into a more compact space of codebooks $\textbf{C}\in\mathbb{R}^{k\times d_{c}}$, where $k\ll n$. The codebooks can be learned by minimizing various objectives such as K-means clustering. The new representation of $x_{i}$ is then computed as follows (Van Den Oord et al., 2017):
$z(x_{i})=c_{k}\quad\text{where}\quad k=\operatorname*{arg\,min}_{j}\|x_{i}-c_{j}\|_{2}^{2}$ (4)
Building upon this concept, RQ-VAE (Lee et al., 2022) extends VQ to a sequence of codebooks, where each consecutive codebook quantizes the residual error from the previous codebook, i.e., $r_{i}=z_{i}-c_{k}$. This hierarchical approach constructs a multi-level quantized representation, enhancing the overall quantization quality.
4 Self-Supervised Graph Tokenization
4.1 Tokenizer Properties
Our goal is to design a graph tokenizer that can learn to generate tokens that exhibit three key characteristics, which are essential for effective graph representation learning. These characteristics are as follows.
Local Interactions. The learned tokens should encapsulate local interactions, allowing the Transformer to focus on global dependencies. This is analogous to Vision Transformers (ViTs), where the Transformer attends to image patches instead of pixels, enabling efficient learning on abstract tokens (Dosovitskiy et al., 2021; Liu et al., 2021). To achieve a similar effect on graph-structured data, we leverage message-passing GNNs as the foundation of the tokenizer’s encoder, capitalizing on their strong locality inductive bias to effectively capture local interactions in the representation space (Battaglia et al., 2018). Our design accommodates various GNN layer choices without constraints; for simplicity, we opt for the widely used Graph Attention Network (GAT) (Veličković et al., 2018) as our base graph encoder. The representation of node $i$ in layer $l$ is computed as:
$h_{i}^{l}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\textbf{W}h_{j}^{(l-1)}\right),\quad\alpha_{ij}=\frac{\exp\left(\sigma\left(\textbf{W}_{2}\left[\textbf{W}_{1}h_{i}^{(l-1)}\|\textbf{W}_{1}h_{j}^{(l-1)}\right]\right)\right)}{\sum\limits_{k\in\mathcal{N}(i)}\exp\left(\sigma\left(\textbf{W}_{2}\left[\textbf{W}_{1}h_{i}^{(l-1)}\|\textbf{W}_{1}h_{k}^{(l-1)}\right]\right)\right)}$ (5)
where $\sigma$ is a non-linearity, and $\alpha_{ij}$ is the normalized attention score between two connected nodes $i$ and $j$.
Memory Efficiency. The tokens should be compact to facilitate efficient memory usage, enabling the Transformer to perform efficient inference. To achieve this, we introduce a Residual-VQ (RVQ) (Lee et al., 2022) layer to quantize the GNN representations into a sequence of discrete tokens. Quantization not only helps with generalization due to its regularization effect but also significantly reduces memory usage. Using an RVQ with $c$ codebooks (typically $c=\{2,\cdots,8\}$), a graph with feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{x}}$ can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{N\times c}$ and codebook representation of $\textbf{C}\in\mathbb{R}^{c\times K\times d_{c}}$, where $c$ is the number of codebooks (i.e., levels of quantization), $K$ is the codebook size, and $d_{c}$ is the code dimension. To illustrate the benefits of this approach, consider a graph with $10^{6}$ nodes and a feature dimension of 1024 ($\textbf{X}\in\mathbb{R}^{10^{6}\times 1024}$). Using an RVQ with 3 codebooks and a codebook size of 256, this graph can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{10^{6}\times 3}$ plus $\textbf{C}\in\mathbb{R}^{3\times 256\times 1024}$, resulting in a 270-fold reduction in required memory.
Robustness and Generalization. The tokens should be robust and generalizable. To achieve this, we rely on graph self-supervised learning. Self-supervised representations have been shown to be more robust to class imbalance (Liu et al., 2022) and distribution shift (Shi et al., 2023), while also capturing better semantic information (Assran et al., 2023) compared to representations learned through supervised objectives. Moreover, self-supervised graph representations have demonstrated superior performance on downstream tasks compared to representations learned in a fully supervised manner, indicating better generalization capabilities (Hu et al., 2020b; Sun et al., 2020; You et al., 2020; 2021; Hassani & Khasahmadi, 2020; Hou et al., 2022; Veličković et al., 2019; Zhu et al., 2020b; Thakoor et al., 2022). Additionally, multi-task learning with self-supervised objectives has been shown to achieve better performance on downstream tasks (Doersch & Zisserman, 2017; Ghiasi et al., 2021). To leverage these benefits, we propose training the GNN encoder with three self-supervised objectives. Unlike RQ-VAE (Lee et al., 2022), which uses reconstruction as its primary objective, we employ graph-specific objectives to capture the nuances of both structure and features within the tokens. Specifically, we use Deep Graph Infomax (DGI) (Veličković et al., 2019) and Graph Masked Auto-Encoder 2 (GMAE2) (Hou et al., 2023). DGI is a contrastive method that contrasts local (node) encoding with global (graph or sub-graph) encoding, whereas GMAE2 combines generative and distillation objectives to jointly reconstruct masked features and track teacher representations.
Figure 1: Overview of our proposed Graph Quantization Transformer (GQT) consisting of three main components: (1) a GNN to encode local interactions, (2) vector quantization for compact representation, and (3) generative and contrastive heads for robust representation learning. We also utilize a Transformers encoder to model long-range interactions. We augment the graph with semantic edges (dashed lines) and generate a sequence for each node based on Personalized PageRank scores. We then modulate the tokens through hierarchical encoding and structural gating, and feed them into the Transformer and aggregate the learned representations through an attention module before passing it to the classification head.
4.2 Training
The GNN encoder is trained through gradient descent to minimize a loss function comprising three terms, where $\beta$ is the loss weight:
$\mathcal{L}=\mathcal{L}_{\text{dgi}}+\mathcal{L}_{\text{gmae2}}+\beta\mathcal{L}_{\text{commit}}$ (6)
The first term is the DGI objective, which maximizes mutual information (MI) between node representations and graph (sub-graph) representations, based on the Jensen-Shannon divergence between the joint and product of marginals as follows (Veličković et al., 2019):
$\mathcal{L}_{\text{dgi}}=\mathbb{E}\left(\sum\limits_{v\in g}\log\left(\mathcal{D}\left(h_{v},h_{g}\right)\right)+\sum\limits_{u\in\tilde{g}}\log\left(1-\mathcal{D}\left(\tilde{h}_{u},h_{g}\right)\right)\right)$ (7)
where $h_{u}$ is the representation of node $u$, $h_{g}$ is the patch (graph/sub-graph) representation that the node belongs to, $\mathcal{D}\left(.,.\right)$ is a discriminator computing the probability scores between local and global information, and $\tilde{g}$ is the corrupted version of the original graph providing negative examples. Following (Veličković et al., 2019), we define the discriminator as a bilinear layer $\mathcal{D}\left(h_{u},h_{g}\right)=\sigma\left(h_{u}^{T}\textbf{W}h_{g}\right)$, compute the global representation as a mean of node representations: $h_{g}=\frac{1}{|\mathcal{V}|}\sum\limits_{v\in g}h_{v}$, and define $\tilde{g}$ as a graph with the same structure but randomly shuffled features.
The second term is the GraphMAE2 objective (Hou et al., 2023), which combines the generative loss of GraphMAE (Hou et al., 2022) with the teacher-(noisy)student distillation loss of BGRL (Thakoor et al., 2022). This combination enables the model to avoid overfitting and learn more semantic representations. The GraphMAE2 loss is computed as follows:
$\mathcal{L}_{\text{gmae2}}=\sum\limits_{v\in\tilde{g}}\left(1-\frac{x_{v}^{T}.\tilde{h}_{v}}{\|x_{v}^{T}\|.\|\tilde{h}_{v}\|}\right)^{\gamma}+\lambda\sum\limits_{v\in g}\left(1-\frac{h_{v}^{T}.\tilde{h}_{v}}{\|h_{v}^{T}\|.\|\tilde{h}_{v}\|}\right)^{\gamma}$ (8)
where $\tilde{g}$ is the masked graph, $\tilde{h}_{v}$ is the node representation of a masked node learned by the noisy student, $h_{v}$ is the corresponding node representation learned by the teacher over the original graph, and $\gamma\geq 1$ is a scaling factor. Note that the teacher’s parameters are updated using an exponential moving average (EMA) of the noisy student’s parameters.
The third term is the commitment loss, which encourages the representations to get close to their corresponding codebook embeddings within the RVQ layer. This loss is computed as:
$\mathcal{L}_{\text{commit}}=\frac{1}{|\mathcal{V}|}\sum\limits_{v\in g}||h_{v}-\text{sg}\left[c_{k}\right]||_{2}$ (9)
where sg is the stop-gradient operator, and $c_{k}$ is the representation of the codebook that $h_{v}$ is assigned to (i.e., the centroid or prototype vector). Note that this loss only affects the node representations and does not update the codebooks.
To initialize and update the codebooks, we employ K-Means clustering and EMA with weight decay $\tau\in[0,1]$, respectively. Specifically, the codebooks are updated as follows:
$c_{k}^{t}=\tau c_{k}^{t-1}+(1-\tau)\frac{1}{|\mathcal{V}_{k}|}\sum\limits_{v\in\mathcal{V}_{k}}h_{v}$ (10)
where $\mathcal{V}_{k}$ is the set of nodes assigned to codebook $c_{k}$. This update rule allows the codebooks to adapt to the changing node representations while maintaining stability.
5 Graph Transformer
5.1 Sequence Generation
Once the tokenizer is trained, each node $v\in\mathcal{V}$ is mapped to a set of $c$ tokens: $T_{v}=\left[t^{v}_{1},\cdots,t^{v}_{c}\right]\in\mathbb{N}^{c}$, which compress information about local interactions. To enable the Transformer to capture long-range interactions, the input should consist of a sequence of tokens from nodes that are likely to have long-range dependencies. To facilitate this, we first augment the graph with semantic edges denoted as $\mathcal{E}_{s}$, which are computed as follows:
$\mathcal{E}_{s}=\left\{e_{u,v}\mid\mathop{\mathrm{arg\,topk}}\limits_{u\in\mathcal{V}}\text{sim}\left(f\left(x_{u}\right),f\left(x_{v}\right)\right)\forall v\in\mathcal{V}\right\}$ (11)
where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $x_{u}$ is the feature vector of node $u$, and $f$ is a projection function. We use cosine similarity as the similarity function and principal component analysis (PCA) as the projection function. This semantic edge augmentation effectively creates sparse edges between each node and its k-nearest neighbors in the feature space, enhancing the model’s ability to recognize and utilize significant long-range dependencies.
We then merge the semantic edges with the original graph edges and use Personalized PageRank (PPR) to generate a sequence per node. A PPR vector for a node $u$ captures the relative importance of other nodes with respect to node $u$ by exploring the graph structure through iterative random walks:
$r=\alpha\mathbf{P}r+(1-\alpha)q$ (12)
where $\mathbf{P}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\in\mathbb{R}^{n\times n}$, $q$ is a stochastic personalized vector, $r$ is the stationary distribution of random walks, and $\alpha$ is a damping factor.
Using PPR enriches the sequence with information beyond local interactions, allowing the Transformer to access potential long-range dependencies. We construct the sequence $S_{v}$ for each node $v$ as follows:
$S_{v}=\left[T_{v}\|T_{u}\|_{u\in\mathop{\mathrm{arg\,topk}}\text{PPR}\left(v,\mathcal{E}\cup\mathcal{E}_{s}\right)}\right]$ (13)
where $S_{v}=\left[t^{v}_{1}\cdots t^{v}_{c}\mid t^{u_{1}}_{1}\cdots t^{u_{1}}_{c}\mid\cdots\mid t^{u_{k}}_{1}\cdots t^{u_{k}}_{c}\right]$ is the sequence of sorted integer tokens with length $c\times(k+1)$, based on the PPR scores for node $u$. Note that the computation of semantic edges and PPR sequences is performed only once as a pre-processing step, which reduces computational overhead during training.
5.2 Token Modulation
There are $c\times K$ possible integer tokens in total, where $c$ is the number of codebooks and $K$ is the codebook size. We randomly initialize an embedding matrix $\textbf{X}_{T}\in\mathbb{R}^{c\times K\times d_{x}}$, which is trained end-to-end with the Transformer. To further enrich the token representation, we introduce an additional token to each node by aggregating the embeddings of its assigned codebooks:
$h_{c}^{v}=\sum\limits_{i=1}^{c}\textbf{C}[i,t_{i}^{v}]$ (14)
We found that adding this explicit aggregated token leads to better performance compared to initializing $\textbf{X}_{T}$ with C. The input representation of the sequence for node $v$ is then defined as:
$S_{v}=\left[\textbf{X}_{T}[i,t^{v}_{i}]\operatorname*{\|}_{i=1}^{c}h_{c}^{v}\ \|\ \textbf{X}_{T}[i,t^{u_{1}}_{i}]\operatorname*{\|}_{i=1}^{c}\ h_{c}^{u_{1}}\ \|\cdots\|\textbf{X}_{T}[i,t^{u_{k}}_{i}]\operatorname*{\|}_{i=1}^{c}\ h_{c}^{u_{k}}\right]$ (15)
This representation combines the individual token embeddings with the aggregated codebook embeddings, providing a more comprehensive and nuanced input to the Transformer.
In order to provide the Transformer with the global structural importance scores of the nodes within the sequence with respect to the target node, we introduce a gating mechanism over the input token embeddings as follows:
$S_{v}=S_{v}\odot\text{Softmax}\left(\mathop{\mathrm{topk}}\text{PPR}\left(v,\mathcal{E}\cup\mathcal{E}_{s}\right)\right)$ (16)
where we first apply a softmax function with temperature $\tau=1$ to normalize the PPR scores, and then multiply each node token’s representation by its corresponding normalized score.
We also introduce two trainable positional encodings to the input tokens. The first positional encoding enables the Transformer to distinguish between tokens from different nodes, while the second encoding, referred to as hierarchical encoding, allows the Transformer to recognize the hierarchy level of each token within the codebooks. We randomly initialize the positional encodings $\textbf{PE}\in\mathbb{R}^{(k+1)\times d_{x}}$ and $\textbf{HE}\in\mathbb{R}^{c\times d_{x}}$ and sum them with the encoding of their corresponding token. For example, the final encoding of the token $j$ of the node $i$ within the sequence is computed as: $x=\textbf{X}_{T}[j,t^{u_{i}}_{j}]+\textbf{PE}[i]+\textbf{HE}[j]$. Note that we did not use any structural encoding, such as Laplacian eigenvectors, as our experiments did not show any significant benefits from including them.
5.3 Transformer Encoder &\&& Classification Head
We use $l$ layers of standard Transformer encoder with flash attention (Dao et al., 2022b) to generate contextual representations per token in the sequence: $\textbf{H}^{(l)}\in\mathbb{R}^{(c+1)\times(k+1)\times d_{h}}$. We then aggregate the token representations for $j$-th node in the sequence by summing along the token dimension:
$\textbf{H}_{v_{j}}=\sum\limits_{i=1}^{c+1}\textbf{H}^{(l)}[i,j]\in\mathbb{R}^{(k+1)\times d_{h}}$ (17)
To obtain a single representation for the entire sequence, We further aggregate the representation using a linear attention layer:
$h=\sum\limits_{i=1}^{k+1}\alpha_{i}h_{i}\quad\text{where}\quad\alpha_{i}=\frac{\exp(\textbf{W}h_{i})}{\sum_{j}\exp(\textbf{W}h_{j})}$ (18)
We feed the resulting representation into a fully-connected classifier and train the model end-to-end with cross-entropy loss. Note that during inference, only the Transformer and classifier are utilized, as the tokenizer is pretrained and the sequences are pre-computed. Furthermore, since we only require discrete tokens and codebook embeddings, our approach allows for efficient memory usage, regardless of graph size enable efficient training and inference on large-scale graphs.
6 Experiments
We comprehensively evaluate GQT on both medium-scale and large-scale node classification tasks, encompassing both homophilous and heterophilous settings across 18 datasets. Homophilous graphs are characterized by nodes with similar classes being connected to each other, whereas heterophilous graphs exhibit connections between nodes with different classes. Following the convention of most existing works on GTs, we focus on node classification tasks in our experiments. However, as discussed in Section 3, our model can be easily extended to graph classification and link prediction tasks. For each evaluation scenario, we adhere to the established experimental protocols from previous works to ensure fair comparisons. Detailed descriptions of the datasets are provided in Appendix A and detailed experimental setup and hyperparameters are provided in Appendix B.
6.1 Comparison with State-of-the-Art
Table 1: Mean node classification performance on medium-scale homophilous datasets over five runs.
Dataset CoraFull CiteSeer PubMed Computer Photo CS Physics WikiCS
#Nodes 19,793 3,327 19,717 13,752 7,650 18,333 34,493 11,701
#Edges 126,842 4,522 88,651 491,722 238,163 163,788 495,924 216,123
#Features 8,710 3,703 500 767 745 6,805 8,415 300
#Classes 70 6 3 10 8 15 5 10
Measure Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$
GNN GCN 61.76$\pm$0.14 76.50$\pm$1.36 86.54$\pm$0.12 89.65$\pm$0.52 92.70$\pm$0.20 92.92$\pm$0.12 96.18$\pm$0.07 77.47$\pm$0.85
GAT 64.47$\pm$0.18 76.55$\pm$1.23 86.32$\pm$0.16 90.78$\pm$0.13 93.87$\pm$0.11 93.61$\pm$0.14 96.17$\pm$0.08 76.91$\pm$0.82
APPNP 65.16$\pm$0.28 76.53$\pm$1.16 88.43$\pm$0.15 90.18$\pm$0.17 94.32$\pm$0.14 94.49$\pm$0.07 96.54$\pm$0.07 78.87$\pm$0.11
GPRGNN 67.12$\pm$0.31 77.13$\pm$1.67 89.34$\pm$0.25 89.32$\pm$0.29 94.49$\pm$0.14 95.13$\pm$0.09 96.85$\pm$0.08 78.12$\pm$0.23
GraphSAINT 67.85$\pm$0.21 -- 88.96$\pm$0.16 90.22$\pm$0.15 91.72$\pm$0.13 94.41$\pm$0.09 96.43$\pm$0.05 --
GraphSAGE -- 75.58$\pm$1.33 87.48$\pm$0.38 91.20$\pm$0.29 94.59$\pm$0.14 93.91$\pm$0.13 96.49$\pm$0.06 74.77$\pm$0.95
PPRGo 63.54$\pm$0.25 -- 87.38$\pm$0.11 88.69$\pm$0.21 93.61$\pm$0.12 92.52$\pm$0.15 95.51$\pm$0.08 78.12$\pm$0.23
GRAND+ 71.37$\pm$0.11 -- 88.64$\pm$0.09 88.74$\pm$0.11 94.75$\pm$0.12 93.92$\pm$0.08 96.47$\pm$0.04 --
GT GT 61.05$\pm$0.38 -- 88.79$\pm$0.12 91.18$\pm$0.17 94.74$\pm$0.13 94.64$\pm$0.13 97.05$\pm$0.05 --
Graphormer OOM -- OOM OOM 92.74$\pm$0.14 94.64$\pm$0.13 OOM --
SAN 59.01$\pm$0.34 -- 88.22$\pm$0.15 89.93$\pm$0.16 94.86$\pm$0.10 94.51$\pm$0.15 OOM --
GraphGPS 55.76$\pm$0.23 76.99$\pm$1.12 88.94$\pm$0.16 OOM 95.06$\pm$0.13 93.93$\pm$0.15 OOM 78.66$\pm$0.49
GOAT -- 76.89$\pm$1.19 86.87$\pm$0.24 90.96$\pm$0.90 92.96$\pm$1.48 94.21$\pm$0.38 96.24$\pm$0.24 77.00$\pm$0.77
NodeFormer -- 76.33$\pm$0.59 89.32$\pm$0.25 86.98$\pm$0.62 93.46$\pm$0.35 95.64$\pm$0.22 96.45$\pm$0.28 74.73$\pm$0.94
DIFFormer -- 76.72$\pm$0.68 89.51$\pm$0.67 91.99$\pm$0.76 95.10$\pm$0.47 94.78$\pm$0.20 96.60$\pm$0.18 73.46$\pm$0.56
NAGphormer 71.51$\pm$0.13 77.42$\pm$1.41 89.70$\pm$0.19 91.22$\pm$0.14 95.49$\pm$0.11 95.75$\pm$0.09 97.34$\pm$0.03 77.16$\pm$0.72
Exphormer 69.09$\pm$0.72 76.83$\pm$1.24 89.52$\pm$0.54 91.59$\pm$0.31 95.27$\pm$0.42 95.77$\pm$0.15 97.16$\pm$0.13 78.54$\pm$0.49
VCR-Graphormer 71.67$\pm$0.10 -- 89.77$\pm$0.15 91.75$\pm$0.15 95.53$\pm$0.14 95.37$\pm$0.04 97.34$\pm$0.04 --
GQT (ours) 71.81$\pm$0.21 77.84$\pm$0.94 90.14$\pm$0.16 92.05$\pm$0.16 95.35$\pm$0.18 96.11$\pm$0.09 97.53$\pm$0.06 79.65$\pm$0.52
Homophilous Node Classification. To evaluate the performance on medium-scale homophilous graphs, we use eight benchmark datasets including CoraFull (Bojchevski & Günnemann, 2017), CiteSeer, and PubMed (Yang et al., 2016), Amazon Computers, Amazon Photos, Co-author CS, and Co-author Physics (Shchur et al., 2018), as well as WikiCS (Mernyei & Cangea, 2020). We compare our results with a comprehensive set of baselines, including four traditional GNNs: GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2019), APPNP (Gasteiger et al., 2018), and GPRGNN (Chien et al., 2020); four scalable GNN variants including GraphSAINT (Zeng et al., 2019), GraphSAGE (Hamilton et al., 2017b), PPRGo (Bojchevski et al., 2020), and GTAND+(Feng et al., 2022); four standard GTs including GT (Dwivedi & Bresson, 2020), Graphormer (Ying et al., 2021b), SAN (Kreuzer et al., 2021), and GraphGPS (Rampášek et al., 2022); and six state-of-the-art scalable GTs including GOAT (Kong et al., 2023a), NodeFormer (Wu et al., 2022a), DiffFormer (Wu et al., 2023a), NAGphormer (Chen et al., 2023), Exphormer (Shirzad et al., 2023), and VCR-Graphormer (Fu et al., 2024). The baseline performance is reported from existing works (Wu et al., 2023b; Luo et al., 2024a; Fu et al., 2024). As shown in Table 1, GQT outperforms the baseline GNN and GT models on 7 out of 8 benchmarks. Notably, this achievement comes with a significant reduction in memory requirement for node features during Transformer training and inference. For example, on the Physics dataset with 34,493 nodes, we only use $256\times 6$ tokens, i.e., 23-fold memory reduction.
Heterophilous Node Classification. Furthermore, we evaluate GQT on six small or medium-scale heterophilous datasets: Squirrel and Chameleon (Rozemberczki et al., 2021), Questions, Roman-Empire, Amazon-Ratings, and Minesweeper (Platonov et al., 2023b). We compare the performance with seven variants of GNNs including GCN, GraphSAGE, GAT, GPRGNN, H2GCN (Zhu et al., 2020a), CPGNN (Zhu et al., 2021), and GloGNN (Li et al., 2022), and six variants of GTs, including GraphGPS, GOAT, NodeFormer, SGFormer, NAGphormer, and Exphormer. The baseline performance is reported from existing works (Wu et al., 2023b; Luo et al., 2024b; Platonov et al., 2023a; Behrouz & Hashemi, 2024). As shown in Table 2, GQT outperforms the baselines on five out of six datasets. We observe that introducing semantic edges and structural gating mechanisms specifically benefits the heterophilous setting (see Appendix C), as they enable the Transformer to capture long-range dependencies that are not easily accessible through the original graph structure.
Large-scale Node Classification We also evaluate the performance of GQT on four large-scale datasets: ogbn-proteins, ogbn-arxiv, ogbn-products (Hu et al., 2020a), and pokec (Leskovec & Krevl, 2014), the last of which is a heterogeneous dataset. We compare the performance against six GNN variants: LINKX (Lim et al., 2021), SIGN (Frasca et al., 2020), GCN, GAT, GraphSAGE, and GPRGNN; and six GT variants: GraphGPS, GOAT, NodeFormer, NAGphormer, Exphormer, and SGFormer (Wu et al., 2024). We report the baseline performance from existing works (Wu et al., 2023b; Luo et al., 2024a). The results (Table 3) show that GQT outperforms the baseline models on all large-scale benchmarks. This is achieved while significantly reducing the required memory. For instance, on the ogbn-products dataset with 2,449,029 nodes and 100-dimensional node features, GQT requires only 3 codebooks of size 4096 each to represent the tokens, resulting in a remarkable 30-fold reduction in memory usage.
Table 2: Mean node classification performance on heterophilous graphs over five runs.
Dataset Squirrel Chameleon Amazon-Ratings Roman-Empire Minesweeper Questions
#Nodes 5,201 2,277 22,662 24,492 10,000 48,921
#Edges 216,933 36,101 32,927 93,050 39,402 153,540
#Features 2,089 2,325 300 300 7 301
#Classes 5 5 18 5 2 2
[…]
GNN GCN 38.67$\pm$1.84 41.31$\pm$3.05 48.70$\pm$0.63 73.69$\pm$0.74 89.75$\pm$0.52 76.09$\pm$1.27
GraphSAGE 36.09$\pm$1.99 37.77$\pm$4.14 53.63$\pm$0.39 85.74$\pm$0.67 93.51$\pm$0.57 76.44$\pm$0.62
GAT 35.62$\pm$2.06 39.21$\pm$3.08 52.70$\pm$0.62 88.75$\pm$0.41 93.91$\pm$0.35 76.79$\pm$0.71
H2GCN 35.10$\pm$1.15 26.75$\pm$3.64 36.47$\pm$0.23 60.11$\pm$0.52 89.71$\pm$0.31 63.59$\pm$1.46
CPGNN 30.04$\pm$2.03 33.00$\pm$3.15 39.79$\pm$0.77 63.96$\pm$0.62 52.03$\pm$5.46 65.96$\pm$1.95
GPRGNN 38.95$\pm$1.99 39.93$\pm$3.30 44.88$\pm$0.34 64.85$\pm$0.27 86.24$\pm$0.61 55.48$\pm$0.91
GloGNN 35.11$\pm$1.24 25.90$\pm$3.58 36.89$\pm$0.14 59.63$\pm$0.69 51.08$\pm$1.23 65.74$\pm$1.19
GT GraphGPS 39.67$\pm$2.84 40.79$\pm$4.03 53.10$\pm$0.42 82.00$\pm$0.61 90.63$\pm$0.67 71.73$\pm$1.47
NodeFormer 38.52$\pm$1.57 34.73$\pm$4.14 43.86$\pm$0.35 64.49$\pm$0.73 86.71$\pm$0.88 74.27$\pm$1.46
SGFormer 41.80$\pm$2.27 44.93$\pm$3.91 48.01$\pm$0.49 79.10$\pm$0.32 90.89$\pm$0.58 72.15$\pm$1.31
NAGphormer 35.80$\pm$1.33 -- 51.26$\pm$0.72 74.34$\pm$0.77 84.19$\pm$0.66 --
Exphormer 36.04$\pm$1.45 -- 53.51$\pm$0.46 89.03$\pm$0.37 90.74$\pm$0.53 --
GQT(ours) 42.54$\pm$1.37 44.23$\pm$3.05 53.89$\pm$0.36 89.21$\pm$0.43 95.28$\pm$0.44 77.28$\pm$1.36
Table 3: Mean node classification performance on large-scale datasets over five runs.
Dataset ogbn-proteins ogbn-arxiv ogbn-products pokec
#Nodes 132,534 169,343 2,449,029 1,632,803
#Edges 39,561,252 1,166,243 61,859,140 30,622,564
#Features 128 8 100 65
#Classes 40 2 47 2
[…]
GNN GCN 72.51$\pm$0.35 71.74$\pm$0.29 75.64$\pm$0.21 75.45$\pm$0.17
GAT 72.02$\pm$0.44 71.95$\pm$0.36 79.45$\pm$0.59 72.23$\pm$0.18
GPRGNN 75.68$\pm$0.49 71.10$\pm$0.12 79.76$\pm$0.59 72.23$\pm$0.18
LINKX 71.37$\pm$0.58 66.18$\pm$0.33 71.59$\pm$0.71 82.04$\pm$0.07
GraphSAGE 77.68$\pm$0.20 71.49$\pm$0.27 78.29$\pm$0.16 75.63$\pm$0.38
SIGN -- 71.95$\pm$0.11 80.52$\pm$0.16 --
GT GraphGPS 76.83$\pm$0.26 70.97$\pm$0.41 OOM OOM
GOAT 74.18$\pm$0.37 72.41$\pm$0.40 82.00$\pm$0.43 66.37$\pm$0.94
NodeFormer 77.45$\pm$1.15 59.90$\pm$0.42 72.93$\pm$0.13 71.00$\pm$1.30
SGFormer 79.53$\pm$0.38 72.63$\pm$0.13 74.16$\pm$0.31 73.76$\pm$0.24
NAGphormer 73.61$\pm$0.33 70.13$\pm$0.55 73.55$\pm$0.21 76.59$\pm$0.25
Exphormer 74.58$\pm$0.26 72.44$\pm$0.28 OOM OOM
GQT(ours) 82.13$\pm$0.34 73.14$\pm$0.16 82.46$\pm$0.17 83.54$\pm$0.26
6.2 Ablation Study
Effect of Tokenization. We examine the performance of the tokenizer by training a linear model on the representations of the learned tokens without modulation, augmentation, or Transformer (1). As shown in Table 4, within the linear evaluation protocol, the tokenizer shows strong performance, surpassing that of GTs such as GraphGPS and NAGphormer, as well as GNNs like GAT and SIGN (Table 3). This implies that the tokenizer is capable of learning effective token representations. To further investigate the importance of the tokenizer, we exclude it and train the Transformer directly on the original node features (2). As expected, this results in significant degradation in performance, highlighting the crucial role of the tokenizer. Additionally, to study the effects of vector quantization, GraphMAE2, and DGI objectives, we train the model by excluding each component (3-5). The results suggest that the SSL objectives contribute more significantly to the performance compared to vector quantization. This is because the primary purpose of vector quantization is to compress information into discrete tokens, reducing memory requirements. Between GraphMAE2 and DGI, GraphMAE2 yields the highest gain. The is due to its composition of two objectives: masked reconstruction and teacher-(noisy)student distillation. Both of these objectives have been shown to outperform InfoMax objectives on downstream tasks (Hou et al., 2022; Thakoor et al., 2022).
Effect of Modulation. We also investigate the impact of codebook embeddings, positional encoding, and structural gating on the model’s performance (6-8). As shown in Table 4, introducing aggregated codebook embeddings leads to improved downstream performance due to the fact that it provides the Transformer with richer representations of each token. Positional encoding, as observed in other domains, contributes moderately to downstream performance. We also note that introducing structural gating yields moderate improvements in homophilous settings, whereas the gains are significant in heterophilous benchmarks (C). This disparity can be attributed to the ability of structural gating to provide the Transformer with importance scores computed over the global graph structure, which is particularly beneficial in heterophilous scenarios.
Effect of Augmentation. We study the effect of semantic edges on downstream performance (9). The results suggest that augmenting the graph structure with semantic edges yields significant gains. This is because introducing semantic edges allows the Transformer to access semantic information that may not be captured by the original graph structure. Furthermore, when combined with random walks, this also enables the Transformer to attend to long-range dependencies. This is especially important in heterophilous benchmarks, where semantic relationships between nodes are more nuanced.
Table 4: Ablation study on effect of proposed components on the ogbn-arxiv dataset.
Graph Tokenizer Token Modulation Augmentation Model Performance
[…]
Embeddings Encoding Gating Edges Sequence
(1) Linear 71.97
(2) Transformer 70.50
(3) Transformer 72.84
(4) Transformer 71.79
(5) Transformer 72.71
(6) Transformer 71.28
(7) Transformer 72.69
(8) Transformer 73.02
(9) Transformer 72.61
(10) Transformer 73.14
7 Conclusion
We introduced GQT (Graph Quantized Tokenizer) to decouple graph tokenization from Transformer using multi-task graph self-supervised learning. The GQT uses vector quantization to learn hierarchical tokens, resulting in significantly reduced memory requirements and improved generalization. We also introduced structural gating, hierarchical encoding, and semantic edges to further improve the performance. We achieved state-of-the-art performance on 16 out of 18 datasets, including large-scale homophilic and heterophilic datasets, while significantly reducing memory requirements. As future directions, we plan to explore the effectiveness of the GQT in graph generative learning by transitioning to a Transformer decoder. Our research lays the groundwork for further investigation into Graph Foundational Models, where LLMs can project heterogeneous features from diverse datasets into a unified textual representation. Building on this foundation, our GQT model can then convert a large number of nodes across different datasets into an efficient set of tokens.
Appendix
Appendix A Datasets
Here we provide a detailed description of the datasets we used. All datasets are publicly available.
  • CoraFull (Bojchevski & Günnemann, 2017), CiteSeer, and Pubmed (Namata et al., 2012) are citation datasets, where nodes represents documents and edges represent citation link. Labels indicates paper category.
  • Computer and Photo (Shchur et al., 2018) are from Amazon co-purchase graph (McAuley et al., 2015), where nodes represent goods and edges indicate that two goods are frequently bought together. Node features are bag-of-words encoded product reviews, and class labels are given by the product category.
  • CS and Physics (Shchur et al., 2018) are co-authorship graphs based on the Microsoft Academic Graph from the KDD Cup 2016 challenges. Here, nodes are authors, that are connected by an edge if they co-authored a paper; node features represent paper keywords for each author’s papers, and class labels indicate most active fields of study for each author.
  • WikiCS (Mernyei & Cangea, 2020) is derived from Wikipedia, where nodes are Computer Science articles, and edges are based on hyperlinks. Nodes are classified into 10 classes representing different branches of the field.
  • Squirrel and Chameleon (Rozemberczki et al., 2021; Pei et al., 2020) are Wikipedia page-page networks, where nodes represent articles from the English Wikipedia, and edges reflect mutual links between them. The nodes were classified into 5 classes in terms of their average monthly traffic.
  • Amazon-Ratings (Platonov et al., 2023b) is based on the Amazon product co-purchasing data. Nodes are products (books, music CDs, DVDs, VHS video tapes), and edges connect products that are frequently bought together. The task is to predict the average rating given to a product by reviewers.
  • Roman-Empire (Platonov et al., 2023b) is based on the Roman Empire article from English Wikipedia. Each node in the graph corresponds to one (non-unique) word in the text. Thus, the number of nodes in the graph is equal to the length of the article. Two words are connected if these words follow each other in the text, or these words are connected in the dependency tree of the sentence. The class of a node is its syntactic role.
  • Minesweeper (Platonov et al., 2023b) is inspired by the Minesweeper game. The graph is a regular 100x100 grid where each node (cell) is connected to eight neighboring nodes (with the exception of nodes at the edge of the grid, which have fewer neighbors). 20% of the nodes are randomly selected as mines. The task is to predict which nodes are mines. The node features are one-hot-encoded numbers of neighboring mines. However, for randomly selected 50% of the nodes, the features are unknown, which is indicated by a separate binary feature.
  • Questions (Platonov et al., 2023b) is based on data from the question-answering website Yandex Q, where nodes are users, and an edge connects two nodes if one user answered the other user’s question during a one-year time interval (from September 2021 to August 2022). The task is to predict which users remained active on the website, forming a binary classification task.
  • ogbn-proteins (Hu et al., 2020a) is a protein-protein assiciation network, where nodes represent proteins, and edges indicate different types of biologically meaningful associations between proteins, e.g., physical interactions, co-expression or homology. The task is to predict the presence of protein functions in a multi-label binary classification setup.
  • ogbn-arxiv (Hu et al., 2020a) is a citation network between all Computer Science (CS) arXiv papers indexed by MAG (Wang et al., 2020). Each node is an arXiv paper and each directed edge indicates that one paper cites another one. The task is to predict the 40 subject areas of arXiv CS papers, e.g., cs.AI, cs.LG, and cs.OS.
  • ogbn-products (Hu et al., 2020a) is an Amazon product co-purchasing network [^1] http://manikvarma.org/downloads/XC/XMLRepository.html of 2M products. Edges indicate that the products are purchased together. The task is to predict the category of a product.
  • pokec (Leskovec & Krevl, 2014; Lim et al., 2021) is a social network, where nodes are users, and edges represent friendships. The task is to predict the gender of users.
For CoraFull, Pubmed, PubMed, Computer, Photo, CS, and Physics, we follow previous work and use 60%/20%/20% train/valid/test split. For WiKiCS, we follow the official split in Mernyei & Cangea (2020). For Squirrel, Chameleon, Amazon-Ratings, Roman-Empire, Minesweeper, and Questions, we follow the splits in Platonov et al. (2023b). For ogbn-proteins, ogbn-arxiv, and ogbn-papers, we follow the splits in Hu et al. (2020a). And for pokec, we follow the split used in Lim et al. (2021).
Appendix B Experimental Setup
Software and hardware. The implementation of our method is based on PyTorch [^2] https://pytorch.org/, PyG [^3] https://pyg.org/, DGL [^4] https://www.dgl.ai/, and vector-quantize-pytorch package [^5] https://github.com/lucidrains/vector-quantize-pytorch. Most of the datasets can be accessed from PyG and DGL. All the experiments are conducted on one Nvidia A100 GPU.
Hyperparameters and experimental details. As illustrated in Figure 1, our method includes two parts: tokenizer and Transformer. We provide the hyperparameters and experimental details for each parts below.
During the training of graph tokenizer, we use full-graph training for small and medium-scale datasets, and apply sampling for large-scale graphs. We consider different sampling methods including random partitioning which randomly samples nodes within a graph and returns their induced subgraph, neighbor sampling (Hamilton et al., 2017b), GraphSAINT (Zeng et al., 2019), and local clustering used in Hou et al. (2023). For the GNN encoder and decoder, we use GCN or GAT as our backbone and tune the number of layers from {1, 2, 3, 4, 5, 6, 7, 8, 9, 10} and hidden dimensions from {128, 256, 512, 1024}. For the quantizer, we use residual-VQ (RVQ) (Lee et al., 2022) and tune the number of codebooks from {1, 2, 3, 6, 9} and codebook size from {128, 256, 512, 1024}. We set the code dimension to be the hidden dimension of the GNN encoder.
During the training of Transformer, we use KNN to add semantic edges and tune the number of semantic neighbors from {0, 5, 10, 15, 20}. Then we use PPR to generate a sequence of nodes for each target node. We tune the number of PPR neighbors from {0, 5, 10, 20, 30, 50}. For the Transformer model, we use the TransformerEncoder module in PyTorch ad our backbone, and tune the number of layers from{1, 2, 3, 4, 5, 6}, number of heads from {4, 8}, and feedforward dimension from {512, 1024, 2048}.
Appendix C Further Ablation Study
Additionally, we provide ablation study on one of the heterophilous dataset. Results are shown in Table 5. Results show that introducing semantic edges and structural gating mechanisms specifically benefits the heterophilous setting.
Table 5: Ablation study on effect of proposed components on the Minesweeper dataset.
Graph Tokenizer Token Modulation Augmentation Model Performance
[…]
Embeddings Encoding Gating Edges Sequence
(1) Linear 90.11
(2) Transformer 90.65
(3) Transformer 95.33
(4) Transformer 92.86
(5) Transformer 93.85
(6) Transformer 93.12
(7) Transformer 94.89
(8) Transformer 93.97
(9) Transformer 92.45
(10) Transformer 95.28

### v2

@@
-Learning Graph Quantized Tokenizers for Transformers
+Learning Graph Quantized Tokenizers
 Abstract
-Transformers serve as the backbone architectures of Foundational Models, where a domain-specific tokenizer helps them adapt to various domains.
-Graph Transformers (GTs) have recently emerged as a leading model in geometric deep learning, outperforming Graph Neural Networks (GNNs) in various graph learning tasks.
-However, the development of tokenizers for graphs has lagged behind other modalities, with existing approaches relying on heuristics or GNNs co-trained with Transformers.
+Transformers serve as the backbone architectures of Foundational Models, where domain-specific tokenizers allow them to adapt to various domains.
+Graph Transformers (GTs) have recently emerged as leading models in geometric deep learning, outperforming Graph Neural Networks (GNNs) in various graph learning tasks.
+However, the development of tokenizers for graphs has lagged behind other modalities.
 To address this, we introduce GQT (Graph Quantized Tokenizer), which decouples tokenizer training from Transformer training by leveraging multi-task graph self-supervised learning, yielding robust and generalizable graph tokens.
 Furthermore, the GQT utilizes Residual Vector Quantization (RVQ) to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities.
-By combining the GQT with token modulation, a Transformer encoder achieves state-of-the-art performance on 16 out of 18 benchmarks, including large-scale homophilic and heterophilic datasets.
+By combining the GQT with token modulation, a Transformer encoder achieves state-of-the-art performance on 20 out of 22 benchmarks, including large-scale homophilic and heterophilic datasets.
 1 Introduction
-Following the success of Transformers (Vaswani et al., 2017) in natural language processing (Devlin et al., 2019; Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021), Graph Transformers (GTs) (Dwivedi & Bresson, 2020; Ying et al., 2021a; Rampášek et al., 2022; Ma et al., 2023; Shirzad et al., 2023; Kong et al., 2023b; Chen et al., 2023; Wu et al., 2022b) have emerged as strong models in geometric deep learning.
-Unlike message-passing Graph Neural Networks (GNNs), which rely on strong locality inductive biases (Battaglia et al., 2018; Veličković et al., 2018; Hou et al., 2020; Hamilton et al., 2017a; Kipf & Welling, 2017), GTs are inherently more expressive due to their ability to capture long-range interactions between nodes (Ma et al., 2023).
-This is particularly beneficial in heterophilous settings where local alignment does not hold (Fu et al., 2024).
-GTs possess an expressive power at least equivalent to the 2-Weisfeiler-Lehman (WL) isomorphism test (Kim et al., 2022), which is sufficient for most real-world tasks (Zopf, 2022).
-This surpasses the expressive power of message-passing GNNs, which are limited to the 1-WL test (Ying et al., 2021a).
-Furthermore, a Transformer with sufficient attention heads can match or exceed the expressive power of a second-order invariant graph network, outperforming message-passing GNNs (Kim et al., 2022).
-However, both GNNs and Transformers are susceptible to over-smoothing (Li et al., 2018; Zhou et al., 2021; Dovonon et al., 2024).
+Following the success of Transformers (Vaswani et al., 2017) in Natural Language Processing (Devlin et al., 2019; Brown et al., 2020) and Computer Vision (Dosovitskiy et al., 2021), Graph Transformers (GTs) (Dwivedi and Bresson, 2020; Ying et al., 2021; Rampášek et al., 2022; Shirzad et al., 2023; Chen et al., 2023; Wu et al., 2022) have emerged as strong models in geometric deep learning.
+In contrast to message-passing Graph Neural Networks (GNNs), which are inherently constrained by strong locality inductive biases (Battaglia et al., 2018; Veličković et al., 2018; Hou et al., 2020; Hamilton et al., 2017; Kipf and Welling, 2017; Zheng et al., 2024), GTs exhibit greater expressivity due to their capacity to capture long-range interactions between nodes (Ma et al., 2023; Kim et al., 2022; Zopf, 2022).
+This is particularly beneficial in heterophilic settings where local alignment does not hold (Fu et al., 2024).
+This dichotomy highlights a fundamental trade-off between GNNs, which focus on local neighborhood aggregation, and GTs, which employ pairwise attention to model global graph structures.
+A natural question arises: can we synergistically integrate the strengths of both approaches to leverage the complementary benefits of local and global representations?
+Specifically, is it possible to harness the locality-aware representations learned by GNNs to construct discrete tokens, thereby enabling GTs to operate efficiently while still capturing salient graph properties?
 GTs require consideration of both graph structure and features, as nodes with identical features will otherwise be projected into the same representation regardless of their surrounding structures (Hoang et al., 2024).
 There are three general approaches to address this limitation (Hoang et al., 2024): (1) node feature modulation, which involves injecting structural information into the node features; (2) context node sampling, where a sampling strategy is used to construct a sequence over the neighbor nodes; and (3) modifying the architecture of a vanilla Transformer to directly incorporate structural biases.
-Given that Transformers are universal approximators of sequence-to-sequence functions (Yun et al., 2020) and considering the rapid developments in efficient implementation of multi-head attention (MHA) module (Dao et al., 2022a; Liu et al., 2024), which enables longer context sizes of up to million-scale tokens (Reid et al., 2024), a well-designed graph tokenizer can allow a vanilla Transformer model to efficiently process even large-scale graphs.
-Recent studies on applying Large Language Models (LLMs) to graph-related tasks have found that representing graphs through textual descriptions can lead to surprisingly strong performance gains that surpass those of GNNs, suggesting that vanilla Transformers are indeed capable of effectively learning graph structures (Ye et al., 2024; He et al., 2024).
-Nonetheless, LLMs are not inference-efficient, and hence our goal in this paper is to devise a lightweight and efficient graph tokenization strategy that enables vanilla Transformer encoders to learn graph structures effectively, without relying on LLMs.
-Tokenizers typically employ self-supervised objectives to abstract data into a sequence of discrete tokens, enabling Transformers to learn representations across various modalities as a unified stream of data.
-This discretization is achieved through vector quantization techniques (Van Den Oord et al., 2017; Lee et al., 2022), which offer several benefits, including: (1) significantly reduced memory requirements, (2) improved inference efficiency, (3) allowing Transformers to focus on long-range dependencies rather than local information, and (4) the capacity to learn more high-level representations due to a compact latent space (Yuan et al., 2021; Yu et al., 2022).
+Given that Transformers are universal approximators of sequence-to-sequence functions (Yun et al., 2020) and considering the rapid developments in efficient implementation of Multi-Head Attention (MHA) module (Dao et al., 2022; Liu et al., 2024), which enables longer context sizes (Reid et al., 2024), we argue that a well-designed graph tokenizer can allow a vanilla Transformer to efficiently process even large-scale graphs.
+Recent studies on applying Large Language Models (LLMs) to Text-Attributed Graphs (TAGs) have shown surprisingly strong performance gains surpassing those of GNNs, suggesting that vanilla Transformers are indeed capable of effectively learning graph structures (Ye et al., 2024; Xu et al., 2025).
+Nonetheless, LLMs are not efficient at inference time.
+Our goal is to devise a lightweight and efficient graph tokenizater that allows vanilla Transformer encoders to learn graph structures effectively.
+Tokenizers typically employ self-supervised objectives to abstract data into a sequence of discrete tokens, allowing Transformers to learn representations across various modalities as a unified stream of data.
+The discretization is usually achieved through vector quantization techniques (Van Den Oord et al., 2017; Lee et al., 2022), which offer several benefits, including: (1) significantly reduced memory requirements, (2) improved inference efficiency, (3) allowing Transformers to focus on long-range dependencies rather than local information, and (4) the capacity to learn more high-level representations due to a compact latent space (Yuan et al., 2021; Yu et al., 2022a).
 These advantages are particularly important in auto-regressive generative modeling, where quantized tokens allow Transformers to generate high-quality outputs in multiple modalities (Dubey et al., 2024; Lee et al., 2022; Dhariwal et al., 2020; Ramesh et al., 2021; Team, 2024).
@@
 More specifically, our contributions are as follows:
-  • We propose a graph tokenizer that utilizes multi-task graph self-supervised objectives to train a graph encoder, enabling it to fully capture local interactions and allowing the Transformer to focus on long-range dependencies.
+  • We propose a graph tokenizer that uses multi-task graph self-supervised learning to train a graph encoder, enabling it to fully capture local interactions and allowing the Transformer to focus on long-range dependencies.
   • Our approach adapts Residual Vector Quantization (RVQ) within the graph tokenizer to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities.
-  • We introduce a novel combination of semantic edges and random walks to facilitate the Transformer’s access to long-range interactions, and employ hierarchical encoding and gating mechanisms to modulate the tokens and provide informative representations to the Transformer.
-  • Through extensive experiments on both homophilic and heterophilic datasets, including large-scale benchmarks, we demonstrate that our proposed tokenizer enables a Transformer encoder to achieve state-of-the-art performance on 16 out of 18 benchmarks while substantially reducing the memory footprint of the embeddings.
+  • We introduce a novel combination of semantic edges and random walks to facilitate access to long-range interactions, and employ hierarchical encoding and gating mechanisms to modulate the tokens and provide informative representations to the Transformer.
+  • Through extensive experiments on both homophilic and heterophilic datasets, including large-scale and long-range benchmarks, we demonstrate that our tokenizer enables Transformer encoders to achieve state-of-the-art performance on 20 out of 22 benchmarks while substantially reducing the memory footprint of the embeddings.
 2 Related Works
 Graph Transformers (GTs) have shown promising performance on various graph learning tasks, surpassing GNNs on many benchmarks.
-GTs can be broadly categorized into two directions (Hoang et al., 2024; Müller et al., 2024): (1) modifying the vanilla Transformer architecture to incorporate structural inductive biases, or (2) encoding the input graph to make it compatible with the vanilla Transformer design.
-Early examples of the first approach include Graph Attention Network (Veličković et al., 2018), which uses an attention module to compute pairwise node attentions and masks the attention matrix based on connectivity information.
-Subsequent works have replaced the scaled-dot attention module with various structure-aware sparse attention modules (Rampášek et al., 2022; Bo et al., 2023; Ying et al., 2021a; Deng et al., 2024; Wu et al., 2023b; Liu et al., 2023a; Chen et al., 2022; Dwivedi & Bresson, 2020; Shirzad et al., 2023; Ma et al., 2023).
-Examples of the second approach include Graph Memory Network (Khasahmadi et al., 2020), which passes non-linear projections of node features and structural encoding to a Transformer-like model.
-Structural encoding methods, such as Laplacian eigenvectors or Random walk-based encoding (Dwivedi et al., 2022; Ma et al., 2023; Cantürk et al., 2024), allow injecting structural information directly into the node features.
-Another approach involves using GNNs to encode local structure along with node features, followed by passing the representation to vanilla Transformers to capture long-range dependencies.
-(Rong et al., 2020; Wu et al., 2021; Chen et al., 2023; 2022).
-Recent studies leverage LLMs, where graphs are represented through natural language expressions, and an LLM performs graph-related tasks through in-context learning, instruction tuning, or soft prompts (Fatemi et al., 2024; Ye et al., 2024; He et al., 2024).
+Designing GTs can be broadly categorized into two directions (Hoang et al., 2024; Müller et al., 2024): (1) modifying the vanilla Transformer architecture to incorporate structural inductive biases, or (2) encoding the input graph to make it compatible with the vanilla Transformer.
+Early examples of the first approach include Graph Attention Network (GAT) (Veličković et al., 2018), which uses an attention module to compute pairwise node attention and masks the attention matrix based on connectivity information.
+Subsequent works have replaced the scaled-dot attention module with various structure-aware sparse attention modules (Rampášek et al., 2022; Bo et al., 2023; Ying et al., 2021; Deng et al., 2024; Wu et al., 2023b; Liu et al., 2023a; Chen et al., 2022; Dwivedi and Bresson, 2020; Shirzad et al., 2023; Ma et al., 2023).
+Graph Memory Network (GMN) (Khasahmadi et al., 2020) is an example of the second approach, which passes non-linear projections of node features and structural encoding to a Transformer-like model.
+Structural encodings such as Laplacian eigenvectors or Random walk-based encoding (Dwivedi et al., 2022a; Ma et al., 2023; Cantürk et al., 2024), allow injecting structural information directly into the node features.
+Some works use GNNs to encode local structure along with node features into embeddings that are passed to vanilla Transformers to capture long-range dependencies (Rong et al., 2020; Wu et al., 2021; Chen et al., 2023, 2022).
+Recent studies leverage LLMs, where graphs are represented through natural language, and an LLM performs graph-related tasks through in-context learning, instruction-tuning, or soft-prompting (Fatemi et al., 2024; Ye et al., 2024; He et al., 2024).
 For a detailed survey on GTs, see (Müller et al., 2024; Hoang et al., 2024).
 Graph Tokenization provides GTs with rich node tokens that encapsulate both structural and semantic information.
-Various approaches have been proposed to define these node tokens.
 TokenGT (Kim et al., 2022) treats nodes and edges as independent tokens defined by their features, type identifiers, and structural encodings.
-NAGphormer (Chen et al., 2023) represents each node as a set of $L$ tokens, where the $l^{th}$ token is the representation of the node from $l^{th}$ hop aggregation.
-In contrast, GraphiT (Mialon et al., 2021) defines a node token as the concatenation of its feature and representation from a graph convolutional kernel network (GCKN).
+NAGphormer (Chen et al., 2023) represents each node with $L$ tokens, where the $l^{th}$ token is the representation of the node from the $l^{th}$ hop aggregation.
+GraphiT (Mialon et al., 2021) defines a node token as the concatenation of its feature and representation from a graph convolutional kernel network (GCKN).
 VCR-Graphormer (Fu et al., 2024) expands the notion of node tokens to include sequences comprising the node feature and features of semantically and community-related neighboring nodes.
-SGT (Liu et al., 2023b) is a non-parametric tokenizer designed for molecular tasks, which simplifies the tokenization process to a non-parametric graph operator without non-linearity, unlike motif-based tokenizers (Zhang et al., 2021; Jin et al., 2018) or GNN pre-training methods (Xia et al., 2023).
+SGT (Liu et al., 2023c) is a non-parametric tokenizer designed for molecular tasks, which unlike motif-based tokenizers (Zhang et al., 2021; Jin et al., 2018) or GNN pre-training methods (Xia et al., 2023), simplifies the tokenization process to a non-parametric graph operator without non-linearity.
 NodePiece (Galkin et al., 2022) is a knowledge-graph tokenizer that represents a target node as a hash of its top-k closest anchors, their distances, and relational context.
-For a more detailed review see Müller et al.
-(2024).
-While vector quantization (VQ) (Van Den Oord et al., 2017; Lee et al., 2022) has been explored in other modalities, its application in graph learning is limited.
-Notable exceptions include VQGraph (Yang et al., 2024), which employs VQ for graph distillation, and NID (Luo et al., 2024a), which uses VQ to learn discrete node IDs for downstream prediction tasks.
+While Vector Quantization (VQ) has been explored in other modalities (Van Den Oord et al., 2017; Lee et al., 2022; Yu et al., 2022b; Van Kempen et al., 2024; Li et al., 2024), its application in graph learning is limited.
+Notable exceptions include VQ-GNN (Ding et al., 2021), which uses quantized representations combined with a low-rank graph convolution matrix to avoid neighbor explosion problem, VQGraph (Yang et al., 2024), which employs VQ for distilling a GNN into an MLP, and NID (Luo et al., 2024a), which uses VQ to learn discrete node IDs for downstream prediction tasks.
 3 Preliminaries
@@
 Let $\mathcal{G}$ denote the space of graphs.
-A graph $g$ within this space is defined as $g=\left(\mathcal{V},\mathcal{E},\textbf{X},\textbf{E}\right)$ where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
+A graph $g\in\mathcal{G}$ is defined as $\left(\mathcal{V},\mathcal{E},\textbf{X},\textbf{E}\right)$ where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
 $\textbf{X}\in\mathbb{R}^{\left|\mathcal{V}\right|\times d_{x}}$ represents the node features of dimension $d_{x}$, and $\textbf{E}\in\mathbb{R}^{\left|\mathcal{V}\right|\times\left|\mathcal{V}\right|\times d_{e}}$ represents the edge features of dimension $d_{e}$.
 A message-passing GNN takes $g$ as input and learns representations $h^{l}_{v}$ for $v\in\mathcal{V}$ ($h^{0}_{v}=x_{v}$) in each layer $l$ as follows (Gilmer et al., 2017):
-$h^{l}_{v}=f^{l}_{\theta}\left(h^{l-1}_{v},g^{l}_{\phi}\left(\left\{\left(h^{l-1}_{v},h^{l-1}_{u},e_{uv}\right)|u\in\mathcal{N}_{i}(v)\right\}\right)\right)$ (1)
-where $f_{\theta}$ and $g_{\phi}$ are known as update (combine) and message (aggregate) functions, respectively, and $\mathcal{N}_{i}(v)$ denotes the set of immediate neighbors of the node $v$.
-With this representation, we can perform various tasks, including node classification as $\text{MLP}\left(h_{v}\right)$, edge prediction as $\text{MLP}\left(h_{u}\odot h_{v}\right)$, or graph classification as $\text{MLP}\left(\mathcal{R}\left(\left\{h_{u}|u\in\mathcal{V}\right\}\right)\right)$ where $\mathcal{R}$ is a pooling (readout) function.
-Graph Transformers, on the other hand, first use a tokenizer $T_{v}=\mathcal{T}_{\psi}\left(\mathcal{N}(v)\right)$ to map each node $v\in\mathcal{V}$ into a sequence of tokens $T_{v}$ by considering some notion of neighborhood $\mathcal{N}$.
+$h^{l}_{v}=f^{l}_{\theta}\left(h^{l-1}_{v},g^{l}_{\phi}\left(\left\{\left(h^{l-1}_{v},h^{l-1}_{u},e_{uv}\right)|u\in\mathcal{N}_{v}\right\}\right)\right)$ (1)
+where $f_{\theta}$ and $g_{\phi}$ are known as combine and aggregate functions, respectively.
+$\mathcal{N}_{v}$ denotes the set of immediate neighbors of the node $v$.
+Once the node representations are computed, we can perform various tasks including node classification as $\text{MLP}\left(h_{v}\right)$, edge prediction as $\text{MLP}\left(h_{u}\odot h_{v}\right)$, or graph classification as $\text{MLP}\left(\mathcal{R}\left(\left\{h_{u}|u\in\mathcal{V}\right\}\right)\right)$, where $\mathcal{R}$ is a pooling (readout) function.
+Graph Transformers use a tokenizer $T_{v}=\mathcal{T}_{\psi}\left(\mathcal{N}(v)\right)$ to map each node $v\in\mathcal{V}$ into a sequence of tokens $T_{v}$ by considering a notion of neighborhood $\mathcal{N}$.
 The simplest design is when $\mathcal{N}$ is zero-hop neighborhood (i.e., the node itself) and $\mathcal{T}_{\psi}$ is a node feature lookup function.
-The neighborhood $\mathcal{N}$ can be extended to include an ego network (Zhao et al., 2021) or top-k Random Walk based neighbors (Fu et al., 2024), and $\mathcal{T}_{\psi}$ can be enhanced to representations from a GNN (Chen et al., 2023).
-Once the tokens are computed, along with a node positional encoding function (PE), we can define the input to a Transformer as $h^{0}_{v}=\left[T_{v}||\text{PE}\left(v\right)\right]$ and compute the representation in each layer $l$ of a vanilla Transformer encoder as follows:
+The neighborhood $\mathcal{N}$ can be extended to include the node’s ego network (Zhao et al., 2021) or top-k Random Walk based neighbors (Fu et al., 2024), and $\mathcal{T}_{\psi}$ can be enhanced to representations from a GNN (Chen et al., 2023).
+Node tokens along with positional encodings (PE) are passed to the Transformer as $h^{0}_{v}=\left[T_{v}||\text{PE}\left(v\right)\right]$.
+The representations in the $l^{th}$ layer of a Transformer encoder are computed as:
 $h^{l}_{v}$ $=\text{LN}\left(\text{MHA}\left(\text{LN}\left(h^{l-1}_{v}\right)\right)+h^{l-1}_{v}\right)$ (2)
@@
 The codebooks can be learned by minimizing various objectives such as K-means clustering.
-The new representation of $x_{i}$ is then computed as follows (Van Den Oord et al., 2017):
-$z(x_{i})=c_{k}\quad\text{where}\quad k=\operatorname*{arg\,min}_{j}\|x_{i}-c_{j}\|_{2}^{2}$ (4)
-Building upon this concept, RQ-VAE (Lee et al., 2022) extends VQ to a sequence of codebooks, where each consecutive codebook quantizes the residual error from the previous codebook, i.e., $r_{i}=z_{i}-c_{k}$.
+The new representation of $x_{i}$ is then computed as (Van Den Oord et al., 2017):
+[…]
+Building upon this concept, Residual-VQ (RVQ) (Lee et al., 2022) extends VQ to a sequence of codebooks, where each consecutive codebook quantizes the residual error from the previous codebook, i.e., $r_{i}=z_{i}-c_{k}$.
 This hierarchical approach constructs a multi-level quantized representation, enhancing the overall quantization quality.
+More details of RVQ are included in Appendix 9.
 4 Self-Supervised Graph Tokenization
 4.1 Tokenizer Properties
-Our goal is to design a graph tokenizer that can learn to generate tokens that exhibit three key characteristics, which are essential for effective graph representation learning.
-These characteristics are as follows.
-Local Interactions.
-The learned tokens should encapsulate local interactions, allowing the Transformer to focus on global dependencies.
-This is analogous to Vision Transformers (ViTs), where the Transformer attends to image patches instead of pixels, enabling efficient learning on abstract tokens (Dosovitskiy et al., 2021; Liu et al., 2021).
-To achieve a similar effect on graph-structured data, we leverage message-passing GNNs as the foundation of the tokenizer’s encoder, capitalizing on their strong locality inductive bias to effectively capture local interactions in the representation space (Battaglia et al., 2018).
-Our design accommodates various GNN layer choices without constraints; for simplicity, we opt for the widely used Graph Attention Network (GAT) (Veličković et al., 2018) as our base graph encoder.
-The representation of node $i$ in layer $l$ is computed as:
-$h_{i}^{l}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\textbf{W}h_{j}^{(l-1)}\right),\quad\alpha_{ij}=\frac{\exp\left(\sigma\left(\textbf{W}_{2}\left[\textbf{W}_{1}h_{i}^{(l-1)}\|\textbf{W}_{1}h_{j}^{(l-1)}\right]\right)\right)}{\sum\limits_{k\in\mathcal{N}(i)}\exp\left(\sigma\left(\textbf{W}_{2}\left[\textbf{W}_{1}h_{i}^{(l-1)}\|\textbf{W}_{1}h_{k}^{(l-1)}\right]\right)\right)}$ (5)
-where $\sigma$ is a non-linearity, and $\alpha_{ij}$ is the normalized attention score between two connected nodes $i$ and $j$.
+Our goal is to design a graph tokenizer that learns node tokens that exhibit three key characteristics:
+Modeling Local Interactions.
+The tokens should encapsulate local interactions, allowing the Transformer to focus on long-range dependencies.
+This is analogous to Vision Transformers (ViTs), where the Transformer attends to image patches instead of pixels (Dosovitskiy et al., 2021; Liu et al., 2021).
+To achieve this, we leverage GNNs as the tokenizer encoder to model local interactions in the representation space (Battaglia et al., 2018).
+Our design accommodates various GNN layer choices without constraints; for simplicity, we opt for a GAT encoder (Veličković et al., 2018).
 Memory Efficiency.
-The tokens should be compact to facilitate efficient memory usage, enabling the Transformer to perform efficient inference.
+The tokens also should be compact to facilitate efficient memory usage.
 To achieve this, we introduce a Residual-VQ (RVQ) (Lee et al., 2022) layer to quantize the GNN representations into a sequence of discrete tokens.
@@
 To illustrate the benefits of this approach, consider a graph with $10^{6}$ nodes and a feature dimension of 1024 ($\textbf{X}\in\mathbb{R}^{10^{6}\times 1024}$).
-Using an RVQ with 3 codebooks and a codebook size of 256, this graph can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{10^{6}\times 3}$ plus $\textbf{C}\in\mathbb{R}^{3\times 256\times 1024}$, resulting in a 270-fold reduction in required memory.
+Using an RVQ with 3 codebooks and a codebook size of 256, this graph can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{10^{6}\times 3}$ plus $\textbf{C}\in\mathbb{R}^{3\times 256\times 1024}$, resulting in a 270-fold reduction in memory.
 Robustness and Generalization.
@@
 Self-supervised representations have been shown to be more robust to class imbalance (Liu et al., 2022) and distribution shift (Shi et al., 2023), while also capturing better semantic information (Assran et al., 2023) compared to representations learned through supervised objectives.
-Moreover, self-supervised graph representations have demonstrated superior performance on downstream tasks compared to representations learned in a fully supervised manner, indicating better generalization capabilities (Hu et al., 2020b; Sun et al., 2020; You et al., 2020; 2021; Hassani & Khasahmadi, 2020; Hou et al., 2022; Veličković et al., 2019; Zhu et al., 2020b; Thakoor et al., 2022).
-Additionally, multi-task learning with self-supervised objectives has been shown to achieve better performance on downstream tasks (Doersch & Zisserman, 2017; Ghiasi et al., 2021).
+Moreover, self-supervised graph representations have demonstrated superior performance on downstream tasks compared to representations learned in a fully supervised manner, indicating better generalization capabilities (Hu et al., 2020b; Sun et al., 2020; You et al., 2020; Fu et al., 2020; You et al., 2021; Hassani and Khasahmadi, 2020; Veličković et al., 2019; Zhu et al., 2020b).
+Additionally, multi-task learning with self-supervised objectives has been shown to achieve better performance on downstream tasks (Doersch and Zisserman, 2017; Ghiasi et al., 2021).
 To leverage these benefits, we propose training the GNN encoder with three self-supervised objectives.
-Unlike RQ-VAE (Lee et al., 2022), which uses reconstruction as its primary objective, we employ graph-specific objectives to capture the nuances of both structure and features within the tokens.
-Specifically, we use Deep Graph Infomax (DGI) (Veličković et al., 2019) and Graph Masked Auto-Encoder 2 (GMAE2) (Hou et al., 2023).
-DGI is a contrastive method that contrasts local (node) encoding with global (graph or sub-graph) encoding, whereas GMAE2 combines generative and distillation objectives to jointly reconstruct masked features and track teacher representations.
+Unlike RQ-VAE (Lee et al., 2022), which uses reconstruction as its primary objective, we employ graph-specific objectives to capture the nuances of both structure and features.
 Figure 1: Overview of our proposed Graph Quantization Transformer (GQT) consisting of three main components: (1) a GNN to encode local interactions, (2) vector quantization for compact representation, and (3) generative and contrastive heads for robust representation learning.
@@
 4.2 Training
-The GNN encoder is trained through gradient descent to minimize a loss function comprising three terms, where $\beta$ is the loss weight:
+To capture different aspects of information, we employ a multi-task learning framework that leverages three distinct families of graph self-supervised objectives: student-teacher distillation (Thakoor et al., 2022), masked autoencoding (Hou et al., 2022), and Infomax (Veličković et al., 2019).
+We also introduce a commitment loss (Van Den Oord et al., 2017) to enforce alignment between learned node representations and the codebook representations.
+Specifically, the GNN encoder is trained through gradient descent to minimize a loss function comprising of three terms, where $\beta$ is the loss weight:
 $\mathcal{L}=\mathcal{L}_{\text{dgi}}+\mathcal{L}_{\text{gmae2}}+\beta\mathcal{L}_{\text{commit}}$ (5)
-The first term is the DGI objective, which maximizes mutual information (MI) between node representations and graph (sub-graph) representations, based on the Jensen-Shannon divergence between the joint and product of marginals as follows (Veličković et al., 2019):
-$\mathcal{L}_{\text{dgi}}=\mathbb{E}\left(\sum\limits_{v\in g}\log\left(\mathcal{D}\left(h_{v},h_{g}\right)\right)+\sum\limits_{u\in\tilde{g}}\log\left(1-\mathcal{D}\left(\tilde{h}_{u},h_{g}\right)\right)\right)$ (7)
-where $h_{u}$ is the representation of node $u$, $h_{g}$ is the patch (graph/sub-graph) representation that the node belongs to, $\mathcal{D}\left(.,.\right)$ is a discriminator computing the probability scores between local and global information, and $\tilde{g}$ is the corrupted version of the original graph providing negative examples.
-Following (Veličković et al., 2019), we define the discriminator as a bilinear layer $\mathcal{D}\left(h_{u},h_{g}\right)=\sigma\left(h_{u}^{T}\textbf{W}h_{g}\right)$, compute the global representation as a mean of node representations: $h_{g}=\frac{1}{|\mathcal{V}|}\sum\limits_{v\in g}h_{v}$, and define $\tilde{g}$ as a graph with the same structure but randomly shuffled features.
+The first term is the Deep Graph Infomax (DGI) (Veličković et al., 2019) objective, which maximizes mutual information (MI) between node representations and graph (sub-graph) representations, based on the Jensen-Shannon divergence between the joint and product of marginals as follows:
+$\mathcal{L}_{\text{dgi}}=\mathbb{E}\left(\sum\limits_{v\in g}\log\left(\mathcal{D}\left(h_{v},h_{g}\right)\right)+\sum\limits_{u\in\tilde{g}}\log\left(1-\mathcal{D}\left(h_{u},h_{g}\right)\right)\right)$ (6)
+where $h_{u}$ is the representation of node $u$.
+$h_{g}$ is the global (sub-graph/graph) representation, computed as the mean of node representations.
+$\tilde{g}$ is the corrupted version of the original graph, with the same structure but randomly shuffled features, providing negative examples for contrastive learning.
+$\mathcal{D}\left(h_{u},h_{g}\right)=\sigma\left(h_{u}^{T}\textbf{W}h_{g}\right)$ is the discriminator that scores whether a node belongs to the graph, and is defined as a bilinear classifier.
 The second term is the GraphMAE2 objective (Hou et al., 2023), which combines the generative loss of GraphMAE (Hou et al., 2022) with the teacher-(noisy)student distillation loss of BGRL (Thakoor et al., 2022).
@@
 where $\tilde{g}$ is the masked graph, $\tilde{h}_{v}$ is the node representation of a masked node learned by the noisy student, $h_{v}$ is the corresponding node representation learned by the teacher over the original graph, and $\gamma\geq 1$ is a scaling factor.
-Note that the teacher’s parameters are updated using an exponential moving average (EMA) of the noisy student’s parameters.
+The teacher’s parameters are updated using an Exponential Moving Average (EMA) of the noisy student’s parameters.
 The third term is the commitment loss, which encourages the representations to get close to their corresponding codebook embeddings within the RVQ layer. This loss is computed as:
@@
 5 Graph Transformer
-5.1 Sequence Generation
-Once the tokenizer is trained, each node $v\in\mathcal{V}$ is mapped to a set of $c$ tokens: $T_{v}=\left[t^{v}_{1},\cdots,t^{v}_{c}\right]\in\mathbb{N}^{c}$, which compress information about local interactions.
+5.1 Graph Serialization
+Once the tokenizer is trained, each node $v\in\mathcal{V}$ is mapped to $c$ discrete tokens: $T_{v}=\left[t^{v}_{1},\cdots,t^{v}_{c}\right]\in\mathbb{N}^{c}$ (i.e., $T_{v}=\textbf{X}_{Q}[v]$), encoding local interactions of that node.
+We then need to serialize the graph in order to input it to the Transformer.
+Semantic Edges.
 To enable the Transformer to capture long-range interactions, the input should consist of a sequence of tokens from nodes that are likely to have long-range dependencies.
 To facilitate this, we first augment the graph with semantic edges denoted as $\mathcal{E}_{s}$, which are computed as follows:
-$\mathcal{E}_{s}=\left\{e_{u,v}\mid\mathop{\mathrm{arg\,topk}}\limits_{u\in\mathcal{V}}\text{sim}\left(f\left(x_{u}\right),f\left(x_{v}\right)\right)\forall v\in\mathcal{V}\right\}$ (11)
+[…]
 where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $x_{u}$ is the feature vector of node $u$, and $f$ is a projection function.
 We use cosine similarity as the similarity function and Principal Component Analysis (PCA) as the projection function.
-This semantic edge augmentation effectively creates sparse edges between each node and its k-nearest neighbors in the feature space, enhancing the model’s ability to recognize and utilize significant long-range dependencies.
-We then merge the semantic edges with the original graph edges and use Personalized PageRank (PPR) to generate a sequence per node.
-A PPR vector for a node $u$ captures the relative importance of other nodes with respect to node $u$ by exploring the graph structure through iterative random walks:
-$r=\alpha\mathbf{P}r+(1-\alpha)q$ (12)
-where $\mathbf{P}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\in\mathbb{R}^{n\times n}$, $q$ is a stochastic personalized vector, $r$ is the stationary distribution of random walks, and $\alpha$ is a damping factor.
-Using PPR enriches the sequence with information beyond local interactions, allowing the Transformer to access potential long-range dependencies.
+The semantic edge augmentation effectively creates sparse edges between each node and its k-nearest neighbors in the feature space, enhancing the model’s ability to recognize and utilize long-range dependencies.
+Structural Serialization.
+We combine the semantic edges with the original edges and use Personalized PageRank (PPR) to generate a sequence per target node.
+This enriches the sequence with information beyond local interactions, allowing the Transformer to access potential long-range dependencies.
 We construct the sequence $S_{v}$ for each node $v$ as follows:
-$S_{v}=\left[T_{v}\|T_{u}\|_{u\in\mathop{\mathrm{arg\,topk}}\text{PPR}\left(v,\mathcal{E}\cup\mathcal{E}_{s}\right)}\right]$ (13)
-where $S_{v}=\left[t^{v}_{1}\cdots t^{v}_{c}\mid t^{u_{1}}_{1}\cdots t^{u_{1}}_{c}\mid\cdots\mid t^{u_{k}}_{1}\cdots t^{u_{k}}_{c}\right]$ is the sequence of sorted integer tokens with length $c\times(k+1)$, based on the PPR scores for node $u$.
-Note that the computation of semantic edges and PPR sequences is performed only once as a pre-processing step, which reduces computational overhead during training.
+[…]
+where $S_{v}=\left[t^{v}_{1}\cdots t^{v}_{c}\mid t^{u_{1}}_{1}\cdots t^{u_{1}}_{c}\mid\cdots\mid t^{u_{k}}_{1}\cdots t^{u_{k}}_{c}\right]$ is a sequence of length $c\times(k+1)$, comprising discrete tokens that represent the target node $v$, followed by discrete tokens of the top-$k$ relevant nodes to node $v$.
+These relevant nodes are determined based on PPR scores.
+Note that the computation of semantic edges and PPR sequences is performed only once as a pre-processing step, thereby reducing the computational overhead during training.
 5.2 Token Modulation
-There are $c\times K$ possible integer tokens in total, where $c$ is the number of codebooks and $K$ is the codebook size.
-We randomly initialize an embedding matrix $\textbf{X}_{T}\in\mathbb{R}^{c\times K\times d_{x}}$, which is trained end-to-end with the Transformer.
-To further enrich the token representation, we introduce an additional token to each node by aggregating the embeddings of its assigned codebooks:
+Token Embeddings.
+There are $c\times K$ possible discrete tokens, where $c$ is the number of codebooks and $K$ is the codebook size.
+We randomly initialize an embedding matrix $\textbf{X}_{T}\in\mathbb{R}^{c\times K\times d_{x}}$, which is trained end-to-end alongside the Transformer.
+To further enrich the token representations, we introduce an additional token for each node that aggregating the embeddings of its assigned codebooks from the pretrained tokenizer:
 $h_{c}^{v}=\sum\limits_{i=1}^{c}\textbf{C}[i,t_{i}^{v}]$ (12)
-We found that adding this explicit aggregated token leads to better performance compared to initializing $\textbf{X}_{T}$ with C.
+where $\textbf{C}[i,j]$ is the embedding corresponding to index j in the ith codebook.
+We found that adding this explicit aggregated token from the codebook leads to better performance compared to initializing $\textbf{X}_{T}$ directly with C.
 The input representation of the sequence for node $v$ is then defined as:
 $S_{v}=\left[\textbf{X}_{T}[i,t^{v}_{i}]\operatorname*{\|}_{i=1}^{c}h_{c}^{v}\ \|\ \textbf{X}_{T}[i,t^{u_{1}}_{i}]\operatorname*{\|}_{i=1}^{c}\ h_{c}^{u_{1}}\ \|\cdots\|\textbf{X}_{T}[i,t^{u_{k}}_{i}]\operatorname*{\|}_{i=1}^{c}\ h_{c}^{u_{k}}\right]$ (13)
-This representation combines the individual token embeddings with the aggregated codebook embeddings, providing a more comprehensive and nuanced input to the Transformer.
+[…]
+Structural Gating.
 In order to provide the Transformer with the global structural importance scores of the nodes within the sequence with respect to the target node, we introduce a gating mechanism over the input token embeddings as follows:
-$S_{v}=S_{v}\odot\text{Softmax}\left(\mathop{\mathrm{topk}}\text{PPR}\left(v,\mathcal{E}\cup\mathcal{E}_{s}\right)\right)$ (16)
+[…]
 where we first apply a softmax function with temperature $\tau=1$ to normalize the PPR scores, and then multiply each node token’s representation by its corresponding normalized score.
+Positional Encoding.
 We also introduce two trainable positional encodings to the input tokens.
@@
 For example, the final encoding of the token $j$ of the node $i$ within the sequence is computed as: $x=\textbf{X}_{T}[j,t^{u_{i}}_{j}]+\textbf{PE}[i]+\textbf{HE}[j]$.
-Note that we did not use any structural encoding, such as Laplacian eigenvectors, as our experiments did not show any significant benefits from including them.
-5.3 Transformer Encoder &\&& Classification Head
-We use $l$ layers of standard Transformer encoder with flash attention (Dao et al., 2022b) to generate contextual representations per token in the sequence: $\textbf{H}^{(l)}\in\mathbb{R}^{(c+1)\times(k+1)\times d_{h}}$.
+Note that we did not use any structural encoding, such as Laplacian eigenvectors, as we did not observe any significant gains from them.
+5.3 Transformer Encoder
+We use $l$ layers of standard Transformer encoder with flash attention (Dao et al., 2022) to generate contextual representations per token in the sequence: $\textbf{H}^{(l)}\in\mathbb{R}^{(c+1)\times(k+1)\times d_{h}}$.
 We then aggregate the token representations for $j$-th node in the sequence by summing along the token dimension:
@@
 $h=\sum\limits_{i=1}^{k+1}\alpha_{i}h_{i}\quad\text{where}\quad\alpha_{i}=\frac{\exp(\textbf{W}h_{i})}{\sum_{j}\exp(\textbf{W}h_{j})}$ (16)
-We feed the resulting representation into a fully-connected classifier and train the model end-to-end with cross-entropy loss.
+We feed the resulting representation into a fully-connected classifier and train the model end-to-end using cross-entropy loss.
 Note that during inference, only the Transformer and classifier are utilized, as the tokenizer is pretrained and the sequences are pre-computed.
-Furthermore, since we only require discrete tokens and codebook embeddings, our approach allows for efficient memory usage, regardless of graph size enable efficient training and inference on large-scale graphs.
+Furthermore, since we only require discrete tokens and codebook embeddings, our approach enables efficient memory usage, regardless of graph size, allowing for efficient training and inference on large-scale graphs.
 6 Experiments
-We comprehensively evaluate GQT on both medium-scale and large-scale node classification tasks, encompassing both homophilous and heterophilous settings across 18 datasets.
-Homophilous graphs are characterized by nodes with similar classes being connected to each other, whereas heterophilous graphs exhibit connections between nodes with different classes.
-Following the convention of most existing works on GTs, we focus on node classification tasks in our experiments.
-However, as discussed in Section 3, our model can be easily extended to graph classification and link prediction tasks.
-For each evaluation scenario, we adhere to the established experimental protocols from previous works to ensure fair comparisons.
-Detailed descriptions of the datasets are provided in Appendix A and detailed experimental setup and hyperparameters are provided in Appendix B.
+We evaluate GQT on both medium- and large-scale graph learning tasks, encompassing 22 homophilic, heterophilic, and long-range benchmarks.
+We follow the established experimental protocols from previous works to ensure fair comparisons.
+Details of the datasets, experimental setup, and hyperparameters are provided in Appendices 10 and 11, respectively.
 6.1 Comparison with State-of-the-Art
-Table 1: Mean node classification performance on medium-scale homophilous datasets over five runs.
+Long-Range Benchmarks.
+We use four datasets from the Long-Range Graph Benchmark (LRGB) (Dwivedi et al., 2022b), including the Peptides-Func dataset for graph classification with Average Precision (AP) metric, the Peptides-Struct dataset for graph regression with Mean Absolute Error (MAE) metric, the COCO-SP dataset for inductive node classification with macro F1 metric, and the PCQM-Contact for link prediction with Mean Reciprocal Rank (MRR) metric.
+We compare our results to baselines reported in (Wang et al., 2024).
+The results shown in Table 1 suggest that GQT is able to capture long-range dependencies and performs well on various graph prediction tasks.
+Table 1: Mean performance on inductive long-range benchmarks over five runs.
+Task Graph Classification Graph Regression Node Classification Link Prediction
+Dataset Peptides-Func Peptides-Struct COCO-SP PCQM-Contact
+#Graphs 15,535 15,535 123,286 529,434
+Avg. #Nodes 150.94 150.94 476.88 30.14
+Avg. #Edges 307.30 307.30 2,693.67 61.09
+Metric AP ↑ MAE ↓ F1 ↑ MRR ↑
+GCN 0.5930$\pm$0.0023 0.3496$\pm$0.0013 0.0841$\pm$0.0010 0.3234$\pm$0.0006
+Exphormer 0.6258$\pm$0.0092 0.2512$\pm$0.0025 0.3430$\pm$0.0108 0.3587$\pm$0.0025
+GPS 0.6535$\pm$0.0041 0.2500$\pm$0.0005 0.3412$\pm$0.0044 0.3337$\pm$0.0006
+Graph-Mamba 0.6739$\pm$0.0087 0.2478$\pm$0.0016 0.3960$\pm$0.0175 0.3395$\pm$0.0013
+GQT (Ours) 0.6903$\pm$0.0085 0.2452$\pm$0.0018 0.4007$\pm$0.0125 0.3427$\pm$0.0012
+Table 2: Mean node classification accuracy on medium-scale homophilic datasets over five runs.
 Dataset CoraFull CiteSeer PubMed Computer Photo CS Physics WikiCS
@@
 #Classes 70 6 3 10 8 15 5 10
-Measure Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$
 GNN GCN 61.76$\pm$0.14 76.50$\pm$1.36 86.54$\pm$0.12 89.65$\pm$0.52 92.70$\pm$0.20 92.92$\pm$0.12 96.18$\pm$0.07 77.47$\pm$0.85
@@
 VCR-Graphormer 71.67$\pm$0.10 -- 89.77$\pm$0.15 91.75$\pm$0.15 95.53$\pm$0.14 95.37$\pm$0.04 97.34$\pm$0.04 --
-GQT (ours) 71.81$\pm$0.21 77.84$\pm$0.94 90.14$\pm$0.16 92.05$\pm$0.16 95.35$\pm$0.18 96.11$\pm$0.09 97.53$\pm$0.06 79.65$\pm$0.52
-Homophilous Node Classification.
-To evaluate the performance on medium-scale homophilous graphs, we use eight benchmark datasets including CoraFull (Bojchevski & Günnemann, 2017), CiteSeer, and PubMed (Yang et al., 2016), Amazon Computers, Amazon Photos, Co-author CS, and Co-author Physics (Shchur et al., 2018), as well as WikiCS (Mernyei & Cangea, 2020).
-We compare our results with a comprehensive set of baselines, including four traditional GNNs: GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2019), APPNP (Gasteiger et al., 2018), and GPRGNN (Chien et al., 2020); four scalable GNN variants including GraphSAINT (Zeng et al., 2019), GraphSAGE (Hamilton et al., 2017b), PPRGo (Bojchevski et al., 2020), and GTAND+(Feng et al., 2022); four standard GTs including GT (Dwivedi & Bresson, 2020), Graphormer (Ying et al., 2021b), SAN (Kreuzer et al., 2021), and GraphGPS (Rampášek et al., 2022); and six state-of-the-art scalable GTs including GOAT (Kong et al., 2023a), NodeFormer (Wu et al., 2022a), DiffFormer (Wu et al., 2023a), NAGphormer (Chen et al., 2023), Exphormer (Shirzad et al., 2023), and VCR-Graphormer (Fu et al., 2024).
-The baseline performance is reported from existing works (Wu et al., 2023b; Luo et al., 2024a; Fu et al., 2024).
-As shown in Table 1, GQT outperforms the baseline GNN and GT models on 7 out of 8 benchmarks.
-Notably, this achievement comes with a significant reduction in memory requirement for node features during Transformer training and inference.
-For example, on the Physics dataset with 34,493 nodes, we only use $256\times 6$ tokens, i.e., 23-fold memory reduction.
-Heterophilous Node Classification.
-Furthermore, we evaluate GQT on six small or medium-scale heterophilous datasets: Squirrel and Chameleon (Rozemberczki et al., 2021), Questions, Roman-Empire, Amazon-Ratings, and Minesweeper (Platonov et al., 2023b).
-We compare the performance with seven variants of GNNs including GCN, GraphSAGE, GAT, GPRGNN, H2GCN (Zhu et al., 2020a), CPGNN (Zhu et al., 2021), and GloGNN (Li et al., 2022), and six variants of GTs, including GraphGPS, GOAT, NodeFormer, SGFormer, NAGphormer, and Exphormer.
-The baseline performance is reported from existing works (Wu et al., 2023b; Luo et al., 2024b; Platonov et al., 2023a; Behrouz & Hashemi, 2024).
+GQT (ours) 71.81$\pm$0.21 77.84$\pm$0.94 90.14$\pm$0.16 93.37$\pm$0.44 95.73$\pm$0.18 96.11$\pm$0.09 97.53$\pm$0.06 80.14$\pm$0.57
+Homophilic Node Classification.
+We use eight medium-scale homophilic datasets including: CoraFull (Bojchevski and Günnemann, 2017), CiteSeer, PubMed (Yang et al., 2016), Amazon Computers, Amazon Photos, Co-author CS, Co-author Physics (Shchur et al., 2018), and WikiCS (Mernyei and Cangea, 2020).
+We compare our results with eight GNNs including: GCN (Kipf and Welling, 2017), GAT, APPNP (Gasteiger et al., 2018), GPRGNN (Chien et al., 2020), GraphSAINT (Zeng et al., 2020), GraphSAGE (Hamilton et al., 2017), PPRGo (Bojchevski et al., 2020), and GTAND+ (Feng et al., 2022).
+We also compare against ten GTs including GT (Dwivedi and Bresson, 2020), Graphormer (Ying et al., 2021), SAN (Kreuzer et al., 2021), GraphGPS (Rampášek et al., 2022), GOAT (Kong et al., 2023), NodeFormer (Wu et al., 2022), DiffFormer (Wu et al., 2023a), NAGphormer (Chen et al., 2023), Exphormer (Shirzad et al., 2023), and VCR-Graphormer (Fu et al., 2024).
+The baseline performance is reported from (Wu et al., 2023b; Luo et al., 2024a).
+GQT outperforms the baseline GNN and GT models on 7 out of 8 benchmarks (Table 2).
+Notably, this achievement comes with a significant memory reduction.
+For example, on the Physics dataset with 34,493 nodes, we only use $256\times 6$ tokens, i.e., a 23-fold memory reduction.
+Heterophilic Node Classification.
+We also evaluate GQT on six medium-scale heterophilic datasets: Squirrel, Chameleon (Rozemberczki et al., 2021), Questions, Roman-Empire, Amazon-Ratings, and Minesweeper (Platonov et al., 2023).
+We compare the performance with seven GNNs: GCN, GraphSAGE, GAT, GPRGNN, H2GCN (Zhu et al., 2020a), CPGNN (Zhu et al., 2021), and GloGNN (Li et al., 2022), and six GTs: GraphGPS, GOAT, NodeFormer, SGFormer, NAGphormer, and Exphormer.
+The baseline performance is reported from (Wu et al., 2023b; Luo et al., 2024b; Platonov et al., 2023; Behrouz and Hashemi, 2024).
 As shown in Table 3, GQT outperforms the baselines on five out of six datasets.
-We observe that introducing semantic edges and structural gating mechanisms specifically benefits the heterophilous setting (see Appendix C), as they enable the Transformer to capture long-range dependencies that are not easily accessible through the original graph structure.
-Large-scale Node Classification We also evaluate the performance of GQT on four large-scale datasets: ogbn-proteins, ogbn-arxiv, ogbn-products (Hu et al., 2020a), and pokec (Leskovec & Krevl, 2014), the last of which is a heterogeneous dataset.
-We compare the performance against six GNN variants: LINKX (Lim et al., 2021), SIGN (Frasca et al., 2020), GCN, GAT, GraphSAGE, and GPRGNN; and six GT variants: GraphGPS, GOAT, NodeFormer, NAGphormer, Exphormer, and SGFormer (Wu et al., 2024).
-We report the baseline performance from existing works (Wu et al., 2023b; Luo et al., 2024a).
+We observe that introducing semantic edges and structural gating specifically benefits the heterophilic setting (Appendix 12.1), as they enable the Transformer to capture long-range dependencies that are not easily accessible through the original graph structure.
+Large-scale Node Classification We also use four large-scale datasets: ogbn-proteins, ogbn-arxiv, ogbn-products (Hu et al., 2020a), and pokec (heterogeneous) (Leskovec and Krevl, 2014).
+We compare the performance against six GNN: LINKX (Lim et al., 2021), SIGN (Frasca et al., 2020), GCN, GAT, GraphSAGE, and GPRGNN; and six GTs: GraphGPS, GOAT, NodeFormer, NAGphormer, Exphormer, and SGFormer (Wu et al., 2023b).
+We report the baseline performance from (Wu et al., 2023b; Luo et al., 2024a).
 The results (Table 4) show that GQT outperforms the baseline models on all large-scale benchmarks.
-This is achieved while significantly reducing the required memory.
-For instance, on the ogbn-products dataset with 2,449,029 nodes and 100-dimensional node features, GQT requires only 3 codebooks of size 4096 each to represent the tokens, resulting in a remarkable 30-fold reduction in memory usage.
-Table 2: Mean node classification performance on heterophilous graphs over five runs.
+This achievement comes with a significant reduction in required memory.
+For instance, on the ogbn-products dataset with 2,449,029 nodes and 100-dimensional node features, GQT requires only 3 codebooks of size 4096, resulting in a 30-fold memory reduction.
+Table 3: Mean node classification performance on heterophilic graphs over five runs.
 Dataset Squirrel Chameleon Amazon-Ratings Roman-Empire Minesweeper Questions
@@
 Exphormer 36.04$\pm$1.45 -- 53.51$\pm$0.46 89.03$\pm$0.37 90.74$\pm$0.53 --
-GQT(ours) 42.54$\pm$1.37 44.23$\pm$3.05 53.89$\pm$0.36 89.21$\pm$0.43 95.28$\pm$0.44 77.28$\pm$1.36
+GQT(ours) 42.72$\pm$1.69 44.23$\pm$3.05 54.32$\pm$0.41 90.98$\pm$0.24 97.36$\pm$0.35 78.94$\pm$0.86
 Table 4: Mean node classification performance on large-scale datasets over five runs.
@@
 Exphormer 74.58$\pm$0.26 72.44$\pm$0.28 OOM OOM
-GQT(ours) 82.13$\pm$0.34 73.14$\pm$0.16 82.46$\pm$0.17 83.54$\pm$0.26
+GQT(ours) 82.13$\pm$0.34 73.14$\pm$0.16 82.46$\pm$0.17 83.76$\pm$0.24
 6.2 Ablation Study
@@
 This is because the primary purpose of vector quantization is to compress information into discrete tokens, reducing memory requirements.
-Between GraphMAE2 and DGI, GraphMAE2 yields the highest gain.
-The is due to its composition of two objectives: masked reconstruction and teacher-(noisy)student distillation.
+Between GraphMAE2 and DGI, the former yields the highest gain.
+This is due to its composition of two objectives: masked reconstruction and teacher-(noisy)student distillation.
 Both of these objectives have been shown to outperform InfoMax objectives on downstream tasks (Hou et al., 2022; Thakoor et al., 2022).
@@
 We also investigate the impact of codebook embeddings, positional encoding, and structural gating on the model’s performance (6-8).
-As shown in Table 4, introducing aggregated codebook embeddings leads to improved downstream performance due to the fact that it provides the Transformer with richer representations of each token.
+As shown in Table 6, introducing aggregated codebook embeddings leads to improved downstream performance because it provides the Transformer with richer representations of each token.
 Positional encoding, as observed in other domains, contributes moderately to downstream performance.
-We also note that introducing structural gating yields moderate improvements in homophilous settings, whereas the gains are significant in heterophilous benchmarks (C).
-This disparity can be attributed to the ability of structural gating to provide the Transformer with importance scores computed over the global graph structure, which is particularly beneficial in heterophilous scenarios.
+We also note that introducing structural gating yields moderate improvements in homophilic settings, whereas the gains are significant in heterophilic benchmarks (12.1).
+This disparity can be attributed to the ability of structural gating to provide the Transformer with importance scores computed over the global graph structure, which is particularly beneficial in heterophilic scenarios.
 Effect of Augmentation.
@@
 This is because introducing semantic edges allows the Transformer to access semantic information that may not be captured by the original graph structure.
-Furthermore, when combined with random walks, this also enables the Transformer to attend to long-range dependencies.
-This is especially important in heterophilous benchmarks, where semantic relationships between nodes are more nuanced.
+Furthermore, when combined with random walks, this also enables the Transformer to attend to long-range dependencies which is particularly important in heterophilic benchmarks, where semantic relationships between nodes are more nuanced.
+Table 5: Accuracy drop under adversarial attack.
+Attack GR-BCD PR-BCD
+PubMed ogbn-arxiv PubMed ogbn-arxiv
+RQ-VAE 20.40% 14.80% 23.30% 17.20%
+GQT (ours) 15.80% 10.40% 18.10% 11.30%
+Robustness Analysis.
+To measure robustness, we use Greedy Randomized Block Coordinate Descent (GRBCD) and Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021) adversarial attacks to measure the accuracy degradation.
+We compare the GQT with an RQ-VAE (Lee et al., 2022).
+The results (Table 5) show that our tokenizer is more robust to attacks.
+This is because GQT is trained with multi-task self-supervised objectives while RQ-VAE is trained with reconstruction objective.
+More details are provided in Appendix 12.
 Table 6: Ablation study on effect of proposed components on the ogbn-arxiv dataset.
@@
 Embeddings Encoding Gating Edges Sequence
-(1) Linear 71.97
-(2) Transformer 70.50
-(3) Transformer 72.84
-(4) Transformer 71.79
-(5) Transformer 72.71
-(6) Transformer 71.28
-(7) Transformer 72.69
-(8) Transformer 73.02
-(9) Transformer 72.61
-(10) Transformer 73.14
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 7 Conclusion
-We introduced GQT (Graph Quantized Tokenizer) to decouple graph tokenization from Transformer using multi-task graph self-supervised learning.
-The GQT uses vector quantization to learn hierarchical tokens, resulting in significantly reduced memory requirements and improved generalization.
-We also introduced structural gating, hierarchical encoding, and semantic edges to further improve the performance.
-We achieved state-of-the-art performance on 16 out of 18 datasets, including large-scale homophilic and heterophilic datasets, while significantly reducing memory requirements.
-As future directions, we plan to explore the effectiveness of the GQT in graph generative learning by transitioning to a Transformer decoder.
-Our research lays the groundwork for further investigation into Graph Foundational Models, where LLMs can project heterogeneous features from diverse datasets into a unified textual representation.
-Building on this foundation, our GQT model can then convert a large number of nodes across different datasets into an efficient set of tokens.
-Appendix
-Appendix A Datasets
-Here we provide a detailed description of the datasets we used. All datasets are publicly available.
-  • CoraFull (Bojchevski & Günnemann, 2017), CiteSeer, and Pubmed (Namata et al., 2012) are citation datasets, where nodes represents documents and edges represent citation link.
-Labels indicates paper category.
-  • Computer and Photo (Shchur et al., 2018) are from Amazon co-purchase graph (McAuley et al., 2015), where nodes represent goods and edges indicate that two goods are frequently bought together.
-Node features are bag-of-words encoded product reviews, and class labels are given by the product category.
-  • CS and Physics (Shchur et al., 2018) are co-authorship graphs based on the Microsoft Academic Graph from the KDD Cup 2016 challenges.
-Here, nodes are authors, that are connected by an edge if they co-authored a paper; node features represent paper keywords for each author’s papers, and class labels indicate most active fields of study for each author.
-  • WikiCS (Mernyei & Cangea, 2020) is derived from Wikipedia, where nodes are Computer Science articles, and edges are based on hyperlinks.
-Nodes are classified into 10 classes representing different branches of the field.
-  • Squirrel and Chameleon (Rozemberczki et al., 2021; Pei et al., 2020) are Wikipedia page-page networks, where nodes represent articles from the English Wikipedia, and edges reflect mutual links between them.
-The nodes were classified into 5 classes in terms of their average monthly traffic.
-  • Amazon-Ratings (Platonov et al., 2023b) is based on the Amazon product co-purchasing data.
-Nodes are products (books, music CDs, DVDs, VHS video tapes), and edges connect products that are frequently bought together.
-The task is to predict the average rating given to a product by reviewers.
-  • Roman-Empire (Platonov et al., 2023b) is based on the Roman Empire article from English Wikipedia.
-Each node in the graph corresponds to one (non-unique) word in the text.
-Thus, the number of nodes in the graph is equal to the length of the article.
-Two words are connected if these words follow each other in the text, or these words are connected in the dependency tree of the sentence.
-The class of a node is its syntactic role.
-  • Minesweeper (Platonov et al., 2023b) is inspired by the Minesweeper game.
-The graph is a regular 100x100 grid where each node (cell) is connected to eight neighboring nodes (with the exception of nodes at the edge of the grid, which have fewer neighbors).
-20% of the nodes are randomly selected as mines.
-The task is to predict which nodes are mines.
-The node features are one-hot-encoded numbers of neighboring mines.
-However, for randomly selected 50% of the nodes, the features are unknown, which is indicated by a separate binary feature.
-  • Questions (Platonov et al., 2023b) is based on data from the question-answering website Yandex Q, where nodes are users, and an edge connects two nodes if one user answered the other user’s question during a one-year time interval (from September 2021 to August 2022).
-The task is to predict which users remained active on the website, forming a binary classification task.
-  • ogbn-proteins (Hu et al., 2020a) is a protein-protein assiciation network, where nodes represent proteins, and edges indicate different types of biologically meaningful associations between proteins, e.g., physical interactions, co-expression or homology.
-The task is to predict the presence of protein functions in a multi-label binary classification setup.
-  • ogbn-arxiv (Hu et al., 2020a) is a citation network between all Computer Science (CS) arXiv papers indexed by MAG (Wang et al., 2020).
-Each node is an arXiv paper and each directed edge indicates that one paper cites another one.
-The task is to predict the 40 subject areas of arXiv CS papers, e.g., cs.AI, cs.LG, and cs.OS.
-  • ogbn-products (Hu et al., 2020a) is an Amazon product co-purchasing network [^1] http://manikvarma.org/downloads/XC/XMLRepository.html of 2M products.
-Edges indicate that the products are purchased together.
-The task is to predict the category of a product.
-  • pokec (Leskovec & Krevl, 2014; Lim et al., 2021) is a social network, where nodes are users, and edges represent friendships. The task is to predict the gender of users.
-For CoraFull, Pubmed, PubMed, Computer, Photo, CS, and Physics, we follow previous work and use 60%/20%/20% train/valid/test split.
-For WiKiCS, we follow the official split in Mernyei & Cangea (2020).
-For Squirrel, Chameleon, Amazon-Ratings, Roman-Empire, Minesweeper, and Questions, we follow the splits in Platonov et al.
-(2023b).
-For ogbn-proteins, ogbn-arxiv, and ogbn-papers, we follow the splits in Hu et al.
-(2020a).
-And for pokec, we follow the split used in Lim et al.
-(2021).
-Appendix B Experimental Setup
-Software and hardware.
-The implementation of our method is based on PyTorch [^2] https://pytorch.org/, PyG [^3] https://pyg.org/, DGL [^4] https://www.dgl.ai/, and vector-quantize-pytorch package [^5] https://github.com/lucidrains/vector-quantize-pytorch.
-Most of the datasets can be accessed from PyG and DGL.
-All the experiments are conducted on one Nvidia A100 GPU.
-Hyperparameters and experimental details.
-As illustrated in Figure 1, our method includes two parts: tokenizer and Transformer.
-We provide the hyperparameters and experimental details for each parts below.
-During the training of graph tokenizer, we use full-graph training for small and medium-scale datasets, and apply sampling for large-scale graphs.
-We consider different sampling methods including random partitioning which randomly samples nodes within a graph and returns their induced subgraph, neighbor sampling (Hamilton et al., 2017b), GraphSAINT (Zeng et al., 2019), and local clustering used in Hou et al.
-(2023).
-For the GNN encoder and decoder, we use GCN or GAT as our backbone and tune the number of layers from {1, 2, 3, 4, 5, 6, 7, 8, 9, 10} and hidden dimensions from {128, 256, 512, 1024}.
-For the quantizer, we use residual-VQ (RVQ) (Lee et al., 2022) and tune the number of codebooks from {1, 2, 3, 6, 9} and codebook size from {128, 256, 512, 1024}.
-We set the code dimension to be the hidden dimension of the GNN encoder.
-During the training of Transformer, we use KNN to add semantic edges and tune the number of semantic neighbors from {0, 5, 10, 15, 20}.
-Then we use PPR to generate a sequence of nodes for each target node.
-We tune the number of PPR neighbors from {0, 5, 10, 20, 30, 50}.
-For the Transformer model, we use the TransformerEncoder module in PyTorch ad our backbone, and tune the number of layers from{1, 2, 3, 4, 5, 6}, number of heads from {4, 8}, and feedforward dimension from {512, 1024, 2048}.
-Appendix C Further Ablation Study
-Additionally, we provide ablation study on one of the heterophilous dataset.
-Results are shown in Table 5.
-Results show that introducing semantic edges and structural gating mechanisms specifically benefits the heterophilous setting.
-Table 5: Ablation study on effect of proposed components on the Minesweeper dataset.
-Graph Tokenizer Token Modulation Augmentation Model Performance
-[…]
-Embeddings Encoding Gating Edges Sequence
-(1) Linear 90.11
-(2) Transformer 90.65
-(3) Transformer 95.33
-(4) Transformer 92.86
-(5) Transformer 93.85
-(6) Transformer 93.12
-(7) Transformer 94.89
-(8) Transformer 93.97
-(9) Transformer 92.45
-(10) Transformer 95.28
+We introduced the Graph Quantized Tokenizer (GQT) to provide standard Transformer encoders to acces discrete graph tokens that encapsulate local interactions and allow Transformers to attend to long-range dependencies within the graph structure.
+This allows us to seamlessly take advantage of the rapid advances in scaling Transformers.
+We achieved state-of-the-art performance on 20 out of 22 datasets, including large-scale and long-range homophilic and heterophilic datasets.
+As future directions, we plan to explore the potential of GQT in generative graph learning.
+Additionally, we aim to couple GQT with LLMs to provide a shared feature space across various graph datasets, paving the way for true Graph Foundational Models (GFMs) (Liu et al., 2023b; Mao et al., 2024).